### v1

# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform Towards Ambient Intelligent Systems

Abstract.
Realizing consumer-grade drones that are as useful as robot vacuums throughout our homes or personal smartphones in our daily lives requires drones to sense, actuate, and respond to general scenarios that may arise. Towards this vision, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks. RASP consists of a mechanical layer to physically swap sensor modules, an electrical layer to maintain power and communication lines to the sensor/actuator, and a software layer to maintain a common interface between the drone and any sensor module in our platform. Leveraging recent advances in large language and visual language models, we further introduce the architecture, implementation, and real-world deployments of a personal assistant system utilizing RASP. We demonstrate that RASP can enable a diverse range of useful tasks in home, office, lab, and other indoor settings.
1. Introduction
Drones are experiencing increased utilization across numerous industries, including search and rescue, construction, delivery, agriculture, and many more (Wingtra, 2023; JOUAV, 2022; heliguy, 2020; Croptracker, 2019; Tucker, 2015; Johnson et al., 2021; Xia et al., 2023). The global commercial drone market is expected to grow to over 580 billion USD by 2030 (dro, 2023). While commercial drones are making significant impacts across various sectors, consumer drones (such as those from DJI (Shenzhen DJI Sciences and Technologies Ltd., 2022) or Skydio (Skydio, Inc., 2016)), that are more common in everyday life, tend to be confined to hobbyist or entertainment purposes, such as aerial photography and videography. However, we believe that the potential of drones in daily life as a versatile platform for mobile sensing and actuation extends far beyond these limited applications.
Imagine a drone in every home, much like robot vacuum cleaners, or each person having a drone acting as a personal “flying smartphone with interchangeable attachements” that you can direct through a variety of interfaces, such as voice commands. The utility of such a platform depends on the quantity and quality of tasks the drone is equipped to handle, which is heavily correlated with the number of sensing and actuation modalities it can support; the number of tasks, generality of a drone, and usefulness of a drone increases with each new sensing and actuation modality it supports.
We propose RASP, an adaptive, modular, and reconfigurable actuation and sensing platform that allows the same drone to adapt to diverse scenarios by quickly and efficiently switching out sensors and/or actuators depending on the situation. Figure 1 shows an overview of the workflow of RASP. The drone lands, communicates with the base station, and specifies the module to swap in. The base station secures the drone and actuates the conveyor belt and gripper to remove and replace the onboard sensor/actuator with the requested module. The unified module interface (UMI) ensures that the sensor/actuator module is firmly attached, connected, and can be accessed by the drone. Finally, the drone flies off with the new module for its new task.
Figure 1. RASP autonomous payload reconfiguration to execute user specified task
Our approach of modularizing a drone’s sensors and actuators offers a number of advantages compared to a static drone with all sensors and actuators attached. Much like single-lens reflex (DSLR) cameras with interchangeable lenses, 1) modularization allows for the creation of an ecosystem of sensors and actuators. This allows consumers to purchase only the sensors and actuators they need and improve drone reusability, expendability, and sustainability. A static drone may not have all the sensors and actuators you may need or want, which requires you to purchase a completely new drone, rather than just the sensor or actuator. 2) Layering and modularization enables you to evolve the drone, sensors, and actuators independently. New sensors and actuators purchased will still be compatible with your current drone as long as the interface remains the same and vice versa. 3) A drone carrying a single sensor or actuator can be designed much smaller, more agile, less noisy, and more suitable for closed environments, such as indoors.
We believe RASP to be one crucial step towards general drone platforms capable of aiding in many daily activities, much like the role of smartphones. To demonstrate this, we prototype and show how RASP could be integrated into a personal assistant system that also leverages static sensors in the environment (cameras) in conjunction with foundational models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting. For example, a drone could carry a temperature sensor to determine the warmest place to sit, before swapping in a payload module to carry a snack to an occupant. To the best of our knowledge, this deployment is the first end-to-end architectural, mechanical, and software design to autonomously control and reconfigure drones for versatile tasks leveraging foundational models. Our contributions are as follows.
1. We create and implement RASP, a modular and reconfigurable sensing and actuation platform that allows a single drone to adapt to a wide range of scenarios. RASP allows drones to dynamically change their onboard sensing and actuation modules in under $25$ seconds.
2. To realize RASP, we propose a three-layer architecture that physically replaces onboard modules, maintains data connection lines, and enables drones to communicate with sensing and actuation modules through a unified interface.
3. To show the importance of RASP for future ambient intelligent systems, we prototype and integrate RASP into a personal assistant system that takes advantage of static sensors in the environment (cameras), large language and vision language models, and the mobility of a drone to realize a wide range of useful applications in smart home, lab, office, and indoor settings.
2. Related Works
There are many reconfigurable and modular sensing platforms in existence. In addition to the platforms provided by open-source do-it-yourself (DIY) electronic vendors, such as Adafruit (Industries, 2022) and Sparkfun (Electronics, 2022), there are also platforms that operate on even lower resource microcontrollers, without an operating system, that are less flexible in the number of interfaces and configurations they support (Yi et al., 2018).  (Zhao et al., 2022, 2023) are reconfigurable sensing platforms based on the Raspberry Pi that have unified and generic hardware interfaces, allowing sensors to use the same set of connectors even if they are interfaced differently in software (e.g., UART, SPI, or I2C). Several platforms leverage the Berkeley TinyOS operating system (Levis et al., 2005), which allow developers to develop extremely long-lasting applications with great flexibility (Culler et al., 2002; O’Flynn et al., 2005; Edmonds et al., 2005; Harte et al., 2007; Mikhaylov et al., 2013; Mikhaylov and Huttunen, 2014). Despite the prevalence of reconfigurable and modular sensing platforms, there currently does not exist any platform that allow drones to reconfigure and adapt its onboard sensors and actuators automatically.
The idea of a smart home and other smart environments has existed for many decades. There are many smart home systems developed by research groups (Das et al., 2002; Mozer, 1998; Helal et al., 2005) and commercial products from companies like Google (LLC, 2016), Amazon (Amazon.com, 2016), and Apple (Inc., 2020). Most of these leverage static sensors and actuators at specific parts of the home to sense and respond to home or human activity. Though there are a few up and coming products (e.g., Ring Always Home Drone (rin, 2023)) and research works (Xia et al., 2021) that leverage drones for indoor and home scenarios, more research is needed in all areas (e.g., control, localization, perception, understanding) to realize a full drone-based system that can operate autonomously and take advantage of the sensing and compute from static sensors throughout indoor environments. (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.
Language and foundational models have seen a surge in usage and research due to their powerful capabilities in allowing computers to understand and interact using natural human modes of communication in an unprecedented manner. There are a few works that leverage language models as an interface for people to interact with their smart homes (King et al., 2023; Nie et al., 2022b, a) or drones and other robotics (Vemprala et al., 2023). However, to the best of our knowledge there are no works that attempt to realize a smart drone-based system for indoor settings that jointly utilizes the strengths of language models, static infrastructure, and a reconfigurable drone platform to perform a diverse range of useful tasks for occupants.
3. Modular and Reconfigurable Sensing and Actuation Drone Platform
Figure 2. RASP architecture.
In this section, we introduce the key components of RASP, our reconfigurable and modular sensing and actuation platform and the drone system that it is built upon. Figure 2 shows the architecture of RASP. At a high level there are three major conceptual architectural components: the ground station or mechanical layer, electrical connection layer, and the software layer. These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 5), sensor/actuation modules (Figures 3c-d and 4), unified module interface (UMI) (Figure 6), the carrier board (Figure 3b), and the physical drone (Figure 3a).
3.1. Architecture
Mechanical Layer. To enable reconfigurability, there needs to be a mechanism to physically swap out different sensor modules; the mechanical layer is responsible for this. The drone will land on the ground station, which will physically detach the current module onboard the drone and attach the newly requested module.
Electrical Connection Layer. Each sensor and actuation module needs to maintain data, communication, and power to the drone, despite being detachable; The electrical layer maintains these connections. We leverage magnets to accomplish this, maintaining physical connection between the module and the drone during a mission, while still allowing the mechanical layer to easily swap out modules.
Software Layer. In addition to maintaining physical and electrical connections, the platform needs to adapt to new modules on-the-fly even if they are swapped out. The software layer provides standardized protocols and libraries that allow the drone to discover, configure, control, and receive data from attached modules.
3.2. Platform Components
Figure 3. a) Physical drone system, b) Carrier board, c) Several sensor modules, d) Actuation module, e) Gripper for swapping sensor and actuation modules
Drone platform. We designed the drone based on the open-source Crazyflie drone (cra, 2023), as shown in Figure 3. We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, not just in open outdoor spaces. To increase its payload, we replaced its brushed motors (19000 Kv powered by a 1-cell battery) with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery). We also printed a protective guard, not only to protect its propellers, but to protect people and other objects from its propeller blades. The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie drone.
Carrier board. The carrier board, which is attached to the base of the drone, provides the physical data, power, and communication connections between the drone platform and each sensor/actuation module. We implement the carrier board on a lightweight \$15 Raspberry Pi Zero 2W (rpi, 2023), which has connections for one sensor or actuation module.
Total Module Fly Drone Module Cost
Mass Mass Time Power Power
Drone Only 344.7g - 3m47s 195.4W - \$344
PM2.5 411.9g 67.2g 3m16s 226.0W 0.29W \$40
[…]
Alcohol 376.2g 31.5g 3m39s 201.7W 0.75W \$8
Actuator 394.8g 50.1g 3m08s 235.1W 1.22W \$7
Table 1. RASP’s supported sensors/actuators and their mass, power consumption, and cost.
Figure 4. a) 3D structure attached to modules to allow the RASP to easily pick them up, b) and c) are displacement plots from stress analyses; the corresponding lowest factors of safety (FOS) are 1.180 and 4.664.
Sensor, Actuation, and Recharging Modules. RASP comes with a collection of sensor and actuation modules (full list in Table 1). These modules are attached or detached from the drone by the ground station/mechanical layer. The actuation module, shown in Figure 3d, is a container structure with a motor that opens and closes a hatch. Small items (e.g., medication, candy, pet food, etc.) can be loaded into this module for the drone to deliver.
Additionally, we create a module that recharges the battery of the drone, which gets attached and removed just like other sensors and actuators. Unlike other standalone sensors and actuators, the recharging module is connected to the wall-powered ground station and charges the drone’s battery at a rate of 1.7 A. At this rate, the fully depleted drone can be fully charged in 30 minutes. The recharging module is the default module that’s attached to the drone each time the drone finishes a task, returns to the ground station, and has no other pending tasks.
To make it easier for the ground station to pick up a module, we 3D printed and attached a structure shown in Figure 4. We performed force analysis (using SOLIDWORKS SimulationXpress) to select the optimal structure and thickness for the frame that is as light as possible to minimize extra weight while being sturdy enough to avoid being deformed when the gripper holds it. We measured that the grippers applies a 5 N force onto the sides of the frame when it grips the frame and a 10 N at the bottom of the frame when it brings the module down to the conveyor belt. To avoid bending at these forces, our force analysis showed when the frame is 1.5mm thick, it had a factor of safety (FOS) greater than 1 in both cases, indicating that the frame was able to withstand the forces applied. The mass of this frame is 11.97 g, which is negligible when compared to the rest of the drone (344.7 g).
Figure 5. Different landing platforms designed and benchmarked. The average landing time took $7.8$ seconds.
Ground station and automated takeoff and landing. The ground station primarily spans the mechanical layer, consisting of a gripper mechanism and a conveyor belt to position modules to swap in/out. We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023) because it provides enough space for the drone to land, as well as enough mechanical control to remove onboard sensors and attach new sensors onto the drone.
We designed and benchmarked three different grippers with different slots and grooves and measured the offset from the center of the parked drone on the ground station. The optimal design, shown in Figure 3e, helps the module slide into a centered position even if the module is slightly offset because of errors in actuating the conveyor to move the module into place.
The ground station contains two additional components: the communication hub and the landing station. The communication hub is the central processor that controls the swapping process. Drones communicate with this hub prior to landing to specify which module it requires and to initiate the swapping process.
The drone then lands at the landing station before the mechanical layer begins the swapping process. The communication hub guides the drone into position for landing using two cameras facing up on top of the ground station (Figure 5b). We also print and attach two ArUco markers to the bottom side of the drone to easily detect the position of the drone and its orientation. ArUco markers, commonly used for camera pose estimation, are similar to QR codes, but carry less encoded information, which makes them more computationally efficient to detect (Sani and Karimian, 2017). In future work, we plan to leverage more complex computer vision models to automatically determine the position and orientation of the drone without needing to add additional markings.
Removing an attached sensor and ensuring that the new sensor is securely fastened requires the drone to land precisely at the specific location where the interchange occurs. However, precise drone landing and alignment is extremely difficult because nonlinearities in drone stability become more pronounced as the drone moves closer to the ground (ground effect) (Matus-Vargas et al., 2021). Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023). As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module.
We test and benchmark two designs for the landing station, as shown in Figure 5c. One design is a normal funnel (Figure 5c-top). The second design contains grooves to help the drone slot into place (Figure 5c-bottom); the base of the drone also has protruding grooves that match the grooves in the funnel. For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar. However, the grooved design corrects for greater drone misalignments from the platform (higher maximum tolerable orientation and offset errors in Figure 5d). As such, we adopt the grooved design into RASP.
Figure 6. Universal module interface for connecting and disconnecting sensors and actuators, as well as recharging the drone.
Unified Module Interface (UMI). The UMI (Figure 6) spans both the electrical connection layer and the software layer, giving the carrier board on the drone a standardized way to communicate with a variety of sensing and actuation modules. The UMI has the following design features.
Standardized 24-pin magnetic connections: Each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, and the carrier board has another set that matches. The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer if a module needs to be switched. The spring-loaded pins enables electrical connections to the sensors and actuators, while allowing the ground station to charge the drone.
Sensor and actuator drivers: The drone needs to support a variety of different sensors and actuators, each of which may be interfaced with differently. To support this, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC. We also provide developers with a driver template that allows them to easily incorporate new sensors and actuators into the RASP ecosystem.
Plug-and-play controller: Once a new module attaches to the carrier board, the drone needs to be able to start collecting data or using the actuator automatically, without needing to manually run any command. To enable this feature, we include a non-volatile memory chip on each module (M24C02 (STMicroelectronics, 2017) 2Kb I2C bus electrically erasable programmable read-only memory (EEPROM)), which stores the identity of the module. The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver from a repository.
4. Towards a Fully Autonomous Drone-based Personal Assistant.
Figure 7. System architecture of intelligent assistant with RASP.
A drone-based personal assistant is one of the core ambient intelligent systems that we see RASP enabling. For several decades now, the number of smart devices, sensors, and actuators deployed throughout our homes, offices, and environments has been growing exponentially (Forecast, 2021). These static sensing devices can provide a coarse-grained and localized picture of events and conditions within the environment; obtaining a fine-grained high resolution picture would require dense deployments of sensors, covering every square inch of the environment, which is impractical. We believe that adapting RASP, drones, and other mobile robotics into smart environments and ecosystems could enable fine-grained sensing and actuation at the tip of our fingers without the need for dense static deployments.
Figure 7 shows the architecture and workflow of our prototype drone-based personal assistant. Static cameras in the environment provide a high-level and coarse-grained view of the environment; in future work and systems, we envision RASP being compatible with other static sensors, beyond cameras. When a user gives a voice command, the system leverages a visual-language model to analyze the scene, interpret the command, and output a list of potential locations of interest, along with the relevant sensor or actuator. Then, the system actuates the RASP drone with the sensor/actuator to each of the locations to complete the task.
For example, one task or question a person may want to know is “where is the warmest place to sit”? A standard RGB camera would not be able to accurately make this determination without additional sensors (e.g., temperature). The voice command and images are analyzed by a visual-language model, which outputs potential locations where a person could sit (e.g., chairs and couches). Then, the system outfits the RASP drone with a temperature module and actuates the drone to each of these locations to measure temperature. The time series data from the temperature sensor will then be analyzed by another language/foundational model to make the final determination. There are four classes of tasks that our system can accommodate.
Object/Location Identification. This set of tasks requires the system to observe the environment and identify an object or location based on the user’s command. For example, the command “where is my phone” requires the visual-language model to identify potential locations for where the phone could be, before sending a drone with a camera to take a closer look and verify.
Object/Location State. This set of tasks involves learning about the state or condition of a specific object or location. For example, “is my food burning” would require the system to identify food that is being cooked, before sending a drone with a CO2 sensor to determine the state of the food.
Surveillance. These tasks differ from the previous two, as it requires the visual-language model to continuously analyze the scene until a potential event occurs, before sending the drone (e.g., “Let me know if any chemical spills on the table”).
Actuation. Unlike the previous three categories, all of which require the drone to sense the environment, actuation tasks all leverage the actuation module to bring an item (e.g., medicine, candy, food) to a specific location.
4.1. Setup and Networked Camera System
To effectively use a camera network to observe the environment, there generally is a lengthy setup process where you need to specify the location of the cameras and have the floor map of the building to stitch together different views. We adopt this self localizing camera network (Liu et al., 2022) that allows users to semi-randomly place cameras on the ceiling, facing downwards perpendicular to the floor; by leveraging motion and people walking throughout daily life, the camera network automatically self-localizes each camera within the network and generates the walkable floormap of the areas observed by the cameras. In this way, users only need to install cameras and do not need to define a floormap or specify and calibrate based on location. More details of our deployment are in Section 5.
4.2. Visual-Language Models for Interpreting Voice Commands
Figure 8. Summary of prompts used to engineer the LLaVA visual-language model (left) to interpret and identify locations of interest given a voice command from a user, as well as for ChatGPT (right) to analyze time series data from the RASP onboard drone sensor.
Figure 9. Segmentation and clustering to break down scenes into smaller more manageable pieces for LLaVA and DINO. (a) Object masks after applying Segment Anything Model (SAM). Extracted frames after clustering object masks based on K-Means (b), hierarchical clustering (c), and ARCK-Means (d). For ARCK-Means, we constrain the aspect ratio of extracted frames to be between 0.67 and 1.5.
Figure 10. Different clustering methods and thresholding evaluated for segmentation. Prompt 1: Describe the image in detail. Prompt 2: Is there a {object name} in the image?
Identifying Objects of Interest. When a user issues a voice command, we pass this command along with the full image of the environment stitched together from all cameras into the open-source Large Language-and-Vision Assistant (LLaVA) visual-language model (Liu et al., 2023a). For the surveillance category, we pass in two consecutive frames or segments to detect changing events. LLaVA outputs (in text) the names of potential objects in the scene that might of interest, as well as the suitable sensor/actuation module that the drone should use. Figure 8-left shows the series of engineered and task-specific prompts we leverage, which was designed based off of the highly active Brex LLM prompt engineering guide (Brex Inc., 2023). The LLaVA text outputs are input as a prompt into Grounding DINO (Liu et al., 2023b), a state-of-art language model for zero-shot object detection, which then outputs bounding boxes and locations of the specified objects.
Object Segmentation and Clustering. However, we found that LLaVA was not able to reliably detect objects of interest in a scene when we passed in an image or a large space (e.g., a single chair in a large room); the larger the scene, the less detailed and more high-level the responses became. To capture small details within the scene that might be relevant to a voice command, we used the Segment Anything Model (SAM) to extract key objects and smaller areas of the scene (Kirillov et al., 2023). Figure 9a shows an example of the object masks that are output by SAM.
We tried inputting these masked objects into LLaVA and then DINO one-by-one, but there were two issues we noticed. First, many of the masks generated often splits objects between two different masks; as such, DINO would often detect the same object multiple times. Second, the masks are not shaped like a regular image (e.g., a square or rectangle); we saw a large degradation in detection performance if we zero-padded the areas around the masked object into a rectangular shape.
To rectify this, we cluster multiple segments together to ensure potentially important objects are all contained with a single rectangular segment. We benchmark the following clustering methods: K-means (Ahmed et al., 2020), hierarchical clustering (Murtagh and Contreras, 2012), and Aspect Ratio Constrained K-Means (ARCK-Means). For K-means and hierarchical clustering, we cluster the centroids of each mask from SAM into $k$ different clusters. Next, we find the minimum-area rectangle that encompasses all masks of each cluster. These $k$ minimum-area rectangles are then analyzed by LLaVA and DINO.
One of the challenges we found with DINO is that it was not able to identify objects of interest very well if input images are very long or very wide. As such, we create and benchmarked a new clustering method called Aspect Ratio Constrained K-Means (ARCK-Means), which constrains the aspect ratio of the clustered masked between two values $x<y$ to maintain a more square shape. To accomplish this, we check if adding a new segment into the existing cluster (during hierarchical clustering) causes the aspect ratio to fall below or above our constraints (e.g., $<x$ or $>y$); If these constraints are broken, then we do not make the assignment.
Figure 9b-d shows an example of segmentation and clustering. We see that for ARCK-Means the clustered masks generated tend to be more square. Additionally, the cluttered table in the bottom right hand corner is fully encompassed by ARCK-Means, but both K-Means and hierarchical clustering do not capture all of the items on the table in a single segment. These improvements of ARCK-Means over the other clustering methods yields higher performance in detecting objects of interest, as shown in Figure 10. In this case and for the rest of the paper, a successful “recall” means that the Grounding DINO model was able to detect and localize the object interest in any one of the segmented clusters it was given, regardless of any additional detections.
First, we tested two prompts into the LLaVA model and vary the size of the object with respect to the frame of one camera (Figure 10a). The first prompt is more general, asking LLaVA to describe the objects present. The second prompt is more specific, asking if the specific object of interest is present in the scene. We see that the more specific a prompt is, the higher the recall, which is the reason why we used more specific prompting (Figure 8). Second, we see that as the size of the object gets smaller with respect to the image frame, the recall gets smaller, since the signal-to-noise ratio of the object gets smaller.
In Figure 10b, we compare the recall of each clustering method after clustering each scene into 5 segments, and see that ARCK-Means has the highest recall due to improvements in the masks it clusters and the shape of the resulting segments; as such, we adopt ARCK-Means into the final system. Figure 10c shows the run time and recall of ARCK-Means as a function of the number of clusters or segments we split the scene into. We see that the recall levels out at around $90\%$ after $5$ clusters, while taking around $100$ms to run the full pipeline. Adding in more clusters does not yield significant improvements in detection, but significantly increases run time; as such, we segment the scene into five segments in the final system. To generate these plots, we used 47 images of indoor home, office, and lab environments. We took 35 images from the ADE20K scene parsing dataset (Zhou et al., 2017) as well as 12 images from our own deployment. We implement and run the full visual-language model pipeline (SAM, clustering, LLaVA, and DINO) on an Nvidia RTX 3090 GPU server.
Prompting users for more context. Some phrases require more context to properly identify locations, sensors, or actuators. For example “tell me the ‘best’ location to sit”. The word “best” could have many meanings (e.g., warmest, coolest, quietest, etc.). If a user gives a command with a non-specific adjective, such as “best”, the system will prompt the user to clarify and be more specific. In the case of an actuation task, the system aims to deliver its payload to one location. If the visual-language model detects multiple potential locations, it will ask the user to clarify the location, either through voice or a web application that we implemented.
4.3. Drone Actuation and Analysis
Once the visual-language models determine important locations, the system attaches the relevant sensor or actuator (from LLaVA) to the RASP drone before moving the drone to each of these locations. In the case of an actuation task, the drone flies to the specified location, opens the hatch to drop the item of interest, before flying back to the ground station.
In the case of a sensing task, the drone flies to each location while reading sensor values from the attached module. After the drone lands back at the ground station, the time series data is input into the GPT-4 language model (OpenAI, 2023) to analyze and make the final determination. Figure 8-right shows the prompts we used for GPT-4. If the sensing modality is to use a camera to identify an object (e.g., a objection/location ID task); the image from the drone is input back into the Grounding DINO model to make the final determination.
4.4. Drone Navigation
Figure 11. a) 2-D drone localizaton CDF. b) Height localization error when drone is at different heights. c) Example ground truth and localized path of drone. d) Localization error of drone vs. length of mission. The localization error remains relatively constant, even as the time of the mission gets longer, demonstrating that the system is not susceptible to localization error drifts. e) deployment floormap. f) camera module used in ceiling camera network.
We use images from the camera network to guide and control the drone. We attach an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3. To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest. Throughout our deployments in Section 5, we observed an median 2-D localization error of 3.29cm, using our camera network. Moreover, as shown in Figure 11, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches.
Using additional markers on top of the drone and the straight line path to navigate is not ideal for generalizing to other drones and environments that are more winding. The focus of this portion of our work is to demonstrate the utility of RASP in an important application (personal assistant) and not specifically drone detection, navigation, or control. Hence, we leave these aspects to future work.
5. Deployment
We deployed the RASP based personal assistant into an office/lab setting as shown in Figure 11e. We used a network of 4 cameras in a uniform grid. As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people walking and moving through the space.
5.1. Benchmarking
Figure 12. Summary of object and event detection by the LLaVA + DINO visual-language pipeline averaged (a) and broken down by category of sensing tasks (b-d).
Refer to caption Refer to caption
Figure 13. Breakdown of accuracy and false positive detections by sensing task category with and without RASP. We see that leveraging RASP in conjunction with static cameras greatly reduces false detections and improves accuracy because the drone can get a closeup view or sense an important part of the environment that a camera alone cannot (e.g., humidity).
Camera Only Camera + RASP
Scenario Precision Recall F-1 Accuracy Sensor Used Precision Recall F-1 Accuracy
Object / Location Identification
Find Phone 68.85% 84.00% 75.68% 73.00% Drone Cam 100.00% 84.00% 91.30% 92.00%
Find Key 78.05% 80.00% 79.01% 71.67% Drone Cam 100.00% 80.00% 88.89% 86.67%
Sit - Temperature 23.47% 92.00% 37.40% 25.24% Temperature 76.67% 92.00% 83.64% 91.26%
Sit - Humidity 25.81% 88.89% 40.00% 28.00% Humidity 82.76% 88.89% 85.71% 92.00%
Sit - Light 20.62% 83.33% 33.06% 22.12% Light Sensor 95.24% 83.33% 88.89% 95.19%
Average (ID) 43.36% 85.64% 53.03% 44.01% 90.93% 85.64% 87.69% 91.42%
State of Object / Location
Faucet Open 80.00% 97.78% 88.00% 82.86% Humidity 93.62% 97.78% 95.65% 94.29%
Stove Open 79.22% 87.14% 82.99% 77.27% Temperature 96.83% 87.14% 91.73% 90.00%
Average (State) 79.61% 92.46% 85.50% 80.06% 95.22% 92.46% 93.69% 92.14%
Surveillance
Food Burning 50.91% 70.00% 58.95% 51.25% PM 93.33% 70.00% 80.00% 82.50%
Chemical Spill 25.40% 80.00% 38.55% 43.33% Gas (Alcohol) 88.89% 80.00% 84.21% 93.33%
Average (Sur.) 38.15% 75.00% 48.75% 47.29% 91.11% 75.00% 82.11% 87.91%
Average (all) 46.54% 83.17% 55.70% 48.99% 91.93% 84.79% 87.78% 90.80%
Table 2. Summary of end-to-end performance between RASP and camera-only for all sensing tasks.
We ran 2-3 tasks in each of the categories of tasks, as we discuss next. For each task, we issued 70 different trials. The scenarios are described in more detail next.
1) Phone. In this ID task, the user asks “where is my phone”. The system will actuate the drone with a camera module to potential locations to detect the phone.
2) Key. This ID task is similar to the phone task, but instead users are looking for keys.
3) Sit - X. In this series of ID tasks, users ask “where is the best place to sit”, based on some sensing modality (e.g., ‘sit - temp’ is where the user asks for the coolest or warmest place). We artificially increase or lower temperatures at different seats by placing space heaters or fans nearby.
4) Faucet. In this state task, users ask “is my faucet still on”; the system will then actuate the drone with a moisture sensor to detect the presence of large amounts of water leaking.
5) Stove. In this state task, users ask “is my stove still on”; the system will then actuate the drone with a temperature sensor to detect the state of the stove.
6) Food. In this surveillance task, users ask “let me know when my food is burning”. For safety, we simulate burning food by boiling water in a kettle, releasing water vapors. The system will then actuate the drone attached with a particulate matter sensor.
7) Chemical. In this surveillance task, users ask “let me know if any chemicals spill”. We will then knock down and spill a glass of alcohol. The system will then actuate the drone with an alcohol sensor to confirm.
8) Medicine. In this actuation task, the user will ask “please bring me my medicine” and wave his/her arms at a camera above. The system will then attach an actuation module loaded with vitamins on the drone, which will then deliver it to the person.
9) Poison. In this actuation task, the user will direct the drone to “deliver rat poison” to a specific location in the environment. The system will load an actuation module with rat poison pellets (simulated with small snacks) and the drone will dispense them at the specified location.
For non-actuation tasks, Figure 12 shows a breakdown of the visual-language model performance in identifying the locations to send the drone versus the segmentation clustering method (Section 4). We see that the ARCK-Means clustering method yields the highest recall, across all scenarios, due to improvements in maintaining a good aspect ratio and object grouping over other clustering methods.
Figure 13 highlights improvements that RASP brings compared to a purely static camera-based system across all sensing tasks, especially with respect to false positives. A successful or “accurate” trial in this context means that the system was able to correctly identify the correct object or location (object/location ID task), correctly identify the state of the object or location (object/location state task), or correctly identify when a targeted event occurs (surveillance task); any additional points of interest identified are counted as incorrect identifications (false positives). For the non-RASP solution, we run the entire SAM, LLaVA, DINO pipeline, but do not actuate the drone. Because we are often looking for small item(s) and locations in a large scene, the visual-language model pipeline often identifies multiple points of interest (e.g., DINO draws multiple bounding boxes and locations). Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high and the precision becomes low with additional locations identified. However, adding in the RASP-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity). This both reduces false positives and improves overall accuracy. Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall) and precision by incorporating RASP.
Scenario # of user prompts # of executions Execution
per execution per battery Time
Object / Location Identification
Find Phone 1.0 7 44.4s
Find Key 1.0 7 46.0s
Sit - Temperature 2.6 3 84.3s
Sit - Humidity 1.9 3 87.7s
Sit - Light 2.3 4 70.4s
Average (ID) 1.8 4.8 66.6s
State of Object / Location
Faucet Open 1.0 6 49.5s
Stove Open 1.0 5 54.2s
Average (State) 1.0 5.5 51.9s
Surveillance
Food Burning 3.1 4 62.5s
Chemical Spill 3.7 5 55.2s
Average (Sur.) 3.4 4.5 58.9s
Average (all) 2.0 4.9 61.6s
Table 3. System performance metrics across benchmarked commands.
For the two actuation tasks, we observed an average offset of the drone, from where it was supposed to travel to drop its payload, of 9.1cm and 10.3cm for the “poison” and “medicine” tasks, respectively. The offset is on orders of centimeters, meaning the system was able to effectively deliver items to the proper location in most cases.
Table 3 shows statistics about the number of tasks per category that could be performed per full charge of battery, execution time and the number of times the system needed to prompt the user for a more specific description or a more accurate location. We see that the average number of user prompts per command is on average 2. For state tasks, this value averaged just 1 prompt (user’s initial command). However, for surveillance and ID tasks, the system often times needed to ask for more information from the user. For all tasks, the average execution time of the visual-language model is on order of seconds, while actuating the drone and analyzing the sensor data is on order of tens of seconds, which is acceptable latency in all of these scenarios. The object/location identification task had a longer average execution time because these tasks generally require the drone to fly and observe multiple locations for each task.
5.2. Reconfiguring Mid-Mission
Category Success P1 Success P2 Success Total
T1: Actuate + Actuate 20/20 = 1.0 18/20 = 0.90 18/20 = 0.90
T2: Sense + Actuate 17/20 = 0.85 20/20 = 1.0 22/25 = 0.85
T3: Sense + Sense 19/20 = 0.95 18/19 = 0.95 18/20 = 0.90
Table 4. Summary of success rate for tasks that require reconfiguring the drone mid-mission, broken down by success rate of the first leg (P1) and second leg (P2).
Figure 14. Example of identified locations and measurements in our multi-step “sense + sense” task.
The previous section demonstrated how RASP could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting. Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy these multi-layered and more complex tasks.
Common tasks that require RASP to reconfigure the drone mid-mission come in three different flavors: 1) actuate + actuate: multiple individual actuation tasks aggregated into one task (e.g., “bring me my medicine AND a snack”), 2) sense + sense: a task that involves multiple sensing modalities (e.g., “find me the coolest (temperature) place to sit out of the sunlight (light)”, 3) sense + actuate: tasks that involve performing actuation in response to sensing (e.g., “place rat poison (actuation) in dark areas (sensing)”). Table 4 summarizes our results, where we show the success rate of the first task (P1), the second task (P2), and the aggregate. We execute each task $20$ times. We discuss each task in more detail, next:
T1: Actuate + Actuate - “Bring medicine AND snack”. In this task, the user wants two (P1 and P2) items brought to him/her. We see that the success rate of delivering both items is high, just like we observed in Section 5.1. The second item (snack) failed two times (P2) because the swapping failed; the drone landed with a high offset from the center of the landing station (Figure 5d).
T2: Sense + Actuate - “Put poison in the warmest area”. In this task, the user may want to place poison (P2) for rats and bugs in warm areas (P1) where they are likely to congregate (e.g., warm places). We simulate “warm places” boiling water in a kettle in locations visible to cameras. There were three times that a location away from the kettle was chosen (P1). In these instances, our image segmentation approach cut out the area we placed the kettle, so the failure point was from the camera rather than RASP. Improving sensing with foundational models in future work is key to realizing a robust version of this end-to-end system.
T3: Sense + Sense - “What is the most humid and brightest location for placing a plant”. Users may want to know a humid (P1) and bright (P2) place for optimal plant growth. We simulate the “correct” location by placing a lamp and kettle at the desired location. There was one instance where the task needed human intervention (failed) because of unsuccessfully swapping in the second sensor; again, the failure came from landing the drone (P2). The foundational model correctly identified the location we placed the lamp, and the drone flew to these locations with the humidity sensor. However, we stopped the run when the drone landed with a high offset that the funnel-shaped landing station could not realign. Another instance failed because our foundational model pipeline did not identify the location we simulated high brightness and humidity (P1). Figure 14 shows an example of a successful run, displaying points where the foundational model identified to send the drone, as well as humidity and light measurements taken by the drone for each of these locations to make the final location determination. Here, we placed the kettle at location three.
Takeaways. We demonstrated how RASP could be used to reconfigure a drone, mid-mission, for tasks that require multiple sensors and/or actuators in a smart building setting with a decent success rate. However, a truly practical system still needs future improvements in drone landing (causes failed module swaps), methods for analyzing sensor data, and fine-grained control of drones.
5.3. In-the-Wild Deployment
Category # Prompts # Exec Exec Time (s) Success
(vLLM + drone)
ID 1.0 21 0.51 + 51.30 15/21 = 0.71
State 1.0 13 0.13 + 37.20 11/13 = 0.85
Surveil. 1.2 22 0.21 + 43.87 19/22 = 0.87
Actuation 2.1 31 0.27 + 29.50 31/31 = 1.0
Table 5. Summary of tasks during in-the-wild deployments. The number of prompts is the average number of times the user needed to prompt the system. This number is often greater than one because either the user used a non-specific adjective (e.g., “best” rather than “warmest” location) or the system needed to narrow down the potential candidate locations in the case of actuation tasks. Number of executions is the total number of tasks issued during the deployment period.
After benchmarking several tasks per category, we allowed people who occupied this office space (Figure 11e) to freely use the system over the course of $5$ days. Table 5 summarizes the number of events that occurred during this period. A total of 8 people issued 87 commands to the system during this time period.
We see that most of the actions issued throughout the deployment were actuation tasks. Around 90% of these tasks involved bringing the user a snack, which we loaded and manually refilled into actuation modules throughout the deployment. ID tasks that users issued generally fell into two categories: finding an area with the least amount of sunlight (our space has many windows and is susceptible to glare) or finding a lost item (e.g., a wallet or phone). For the surveillance and the state tasks, most users asked the system about a 3D print job, whether a heat element was left on (e.g., soldering iron), or if there were anyone occupying different parts of the space.
The category of tasks that had the lowest success rate was the object/location ID category. This is because most of these tasks relied on static cameras or the camera on the drone to find something extremely small in the landscape of a scene (e.g., a circuit component or a phone), making it difficult for the visual-language pipeline to identify relevant locations. Even after flying the drone to the specified location, it can be difficult to detect; we envision future work focusing on how to design search algorithms and protocols for drones to identify small objects of interest. Several items that users wanted the drone to look for were also underneath furniture or tables; a camera mounted on the ceiling or walls have limited view of these items. Another avenue of research for realizing a drone or robot-based personal assistant could be how to leverage and design small robotic systems (e.g., physical design, path planning, search algorithms, etc.) to reach and look for items in areas unobservable by static sensor deployments. On the flip side, the actuation task had the highest success rate particularly because the system prompted users each time to confirm the location to make the delivery, which reduces reliance on language models and perception algorithms to make this determination. Although there are still improvements needed to realize a truly autonomous drone-based personal assistant, all users were positively receptive to this system and could see its value.
6. Discussion
The focus of this work is to realize an autonomous reconfigurable and modular sensing and actuation platform drone platform that can adapt its sensing and actuation capabilities depending on the task at hand. As such, a single drone can be adapted for multiple tasks. We demonstrate the utility of such a platform for ambient intelligent systems by integrating it with a drone-based personal assistant that we envision will become commonplace in future homes, offices, and buildings. Specifically, we demonstrate the potential of leveraging static sensors in the environment and foundational models (e.g., visual-language models) in conjunction with a powerful and adaptable mobile platform in smart building, home, office, and indoor environments. To realize a fully autonomous drone-based personal assistant and other intelligent systems that can fully take advantage of RASP, there are several directions that need more exploration.
Robust path planning, navigation, and search for indoor environments. As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented on the drone that were not able to capture and detect small objects of interest in a cluttered home/indoor setting. More work into exploration and detecting these small objects of interest is needed.
Penetrative AI, foundational models and leveraging diverse sensors in the environment. In this work, we leveraged cameras in conjunction with language to understand and interpret the environment. However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundational models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing our platform and case study to all types of scenarios.
Drone noise reduction, smaller sizes, and power efficiency. Drones and quadcopters are very loud because of its propellers. Having them comfortably co-exist with people in indoor environments requires more research into the design of them to reduce their size and noise, while prolonging their battery life and enabling automatic charging while docked.
7. Conclusion
We propose RASP, a modular and reconfigurable sensing and actuation platform that allows small drones to swap sensing and actuation capabilities dynamically. RASP allows a single drone to reconfigure itself to satisfy a diverse range of tasks that is commonly not possible with current drones. Current industrial drones are typically designed to satisfy a single application domain, and consumer drones are currently mostly used by hobbyists for entertainment. We demonstrate the diverse range of tasks RASP can enable by integrating RASP into a personal assistant system that also leverages static cameras to enable a wide range of useful tasks that were previously not possible. RASP is a step towards generalizing the use and presence of drones in daily life, much like how smartphones have become ubiquitous.
Acknowledgements.
This research was partially supported by COGNISENSE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA, as well as the National Science Foundation under Grant Number CNS-1943396. The views and conclusions contained here are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Columbia University, NSF, SRC, DARPA, or the U.S. Government or any of its agencies.

### v2

@@
-RASP: A Drone-based Reconfigurable Actuation and Sensing Platform Towards Ambient Intelligent Systems
+RASP: A Drone-based Reconfigurable Actuation and Sensing Platform for Engaging Physical Environments with Foundation Models
 Abstract.
-Realizing consumer-grade drones that are as useful as robot vacuums throughout our homes or personal smartphones in our daily lives requires drones to sense, actuate, and respond to general scenarios that may arise.
-Towards this vision, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks.
-RASP consists of a mechanical layer to physically swap sensor modules, an electrical layer to maintain power and communication lines to the sensor/actuator, and a software layer to maintain a common interface between the drone and any sensor module in our platform.
-Leveraging recent advances in large language and visual language models, we further introduce the architecture, implementation, and real-world deployments of a personal assistant system utilizing RASP.
-We demonstrate that RASP can enable a diverse range of useful tasks in home, office, lab, and other indoor settings.
+Foundation models and large language models have shown immense human-like understanding and capabilities for generating text and digital media.
+However, foundation models that can freely sense, interact, and actuate the physical world like in the digital domain is far from being realized.
+This is due to a number of challenges including: 1) being constrained to the types of static devices and sensors deployed, 2) events often being localized to one part of a large space, and 3) requiring dense and deployments of devices to achieve full coverage.
+As a critical step towards enabling foundation models to successfully and freely interact with the physical environment, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks.
+We demonstrate through real smart home deployments that RASP enables FMs and LLMs to complete diverse tasks up to $85\%$ more successfully by allowing them to target specific areas with specific sensors and actuators on-the-fly.
 1. Introduction
-Drones are experiencing increased utilization across numerous industries, including search and rescue, construction, delivery, agriculture, and many more (Wingtra, 2023; JOUAV, 2022; heliguy, 2020; Croptracker, 2019; Tucker, 2015; Johnson et al., 2021; Xia et al., 2023).
-The global commercial drone market is expected to grow to over 580 billion USD by 2030 (dro, 2023).
-While commercial drones are making significant impacts across various sectors, consumer drones (such as those from DJI (Shenzhen DJI Sciences and Technologies Ltd., 2022) or Skydio (Skydio, Inc., 2016)), that are more common in everyday life, tend to be confined to hobbyist or entertainment purposes, such as aerial photography and videography.
-However, we believe that the potential of drones in daily life as a versatile platform for mobile sensing and actuation extends far beyond these limited applications.
-Imagine a drone in every home, much like robot vacuum cleaners, or each person having a drone acting as a personal “flying smartphone with interchangeable attachements” that you can direct through a variety of interfaces, such as voice commands.
-The utility of such a platform depends on the quantity and quality of tasks the drone is equipped to handle, which is heavily correlated with the number of sensing and actuation modalities it can support; the number of tasks, generality of a drone, and usefulness of a drone increases with each new sensing and actuation modality it supports.
-We propose RASP, an adaptive, modular, and reconfigurable actuation and sensing platform that allows the same drone to adapt to diverse scenarios by quickly and efficiently switching out sensors and/or actuators depending on the situation.
-Figure 1 shows an overview of the workflow of RASP.
+Foundation models (FM) and artificial general intelligence (AGI) have the potential to bring about massive technological and social change to the physical world by discovering, performing, and continually adapting to new applications and environments without needing to manually embed or program them.
+While LLMs and FMs have been shown to have human-like understanding and digital media generation capabilities, their ability to respond to events and actuate systems in the physical world are comparatively less explored.
+There are a number of works that incorporate language models to control physical robotic systems through voice (Vemprala et al., 2023; Nie et al., 2022b), from both industry and academia, such as the Figure01 AGI robot (Figure, 2024).
+The sensing and control in such scenarios are often localized to the vicinity of the autonomous agent.
+Scaling up, there are few works that explore the use of LLMs to actuate our environments, particularly smart homes (King et al., 2023, 2024; Rivkin et al., 2023), where events and actions may occur anywhere in the space.
+These works generally focus on adapting LLMs as a human-like interface to actuate common internet-connected smart appliances (e.g., speakers, television, air conditioning, etc.).
+Much like how FMs enable general human language and sensory understanding and responses, our work explores how FMs could be used to enable a diverse range of general interactions with objects and spaces that likely cannot be actuated digitally through code.
+These interactions form a large portion of our daily interactions.
+Figure 1. RASP autonomous payload reconfiguration to execute user specified task
+For example, a person may want his/her home to physically bring a snack or medicine; this would require a robotic system that could physically identify and carry the payload.
+Another person, coming into an office area for the time may want to know the warmest desk to sit at in the building; this would require the building to utilize a dense deployment of temperature sensors.
+A third person in a chemical lab may want the building to monitor and notify him/her about the results of a chemical reaction that s/he started before leaving; this situation could likely be accomplished with a camera with special capabilities to detect these reactions.
+While it is possible to enable all these applications, these examples highlight two main challenges that prevent FMs and LLMs for achieving the same amount of autonomy in our physical environments as we have seen in the digital domain:
+1.
+Each new application requires a new device or sensor, often with built-in special capabilities.
+Each of the three examples we discussed would require the user to purchase or engineer a system to satisfy that task.
+Evolving new applications and functionality in this way is also not scalable.
+Another way to state this challenge is that the generality and types of applications available are limited by the functionality of the devices and sensors present.
+2.
+Events, tasks, and actions are often localized.
+Certain events and tasks a user may require the FM to analyze the entire space, but perform an action at a target location.
+For example, placing rat poison in areas around the home that will be most effective.
+If a specific sensing modality needed to make this judgment (e.g., moisture) is not present at a specific location, then the FM cannot obtain the information it needs to carry out the task.
+This leads to the third challenge.
+3.
+Achieving full coverage across an entire space requires space-dependent dense deployments.
+The example of finding the “best desk” to sit requires a dense deployment of temperature sensors.
+While it is common for smart spaces to deploy smart devices and sensors, new applications may require dense deployments that need to be tailored to the layout of the environment, making it difficult to achieve full coverage for each new application.
+Towards enabling general task execution by FMs in physical environments, we propose RASP, an adaptive, modular, and reconfigurable actuation and sensing drone platform that allows FMs and LLMs to adapt the same drone to diverse scenarios by efficiently switching out sensors and/or actuators depending on the situation.
+Figure 1 shows an overview of RASP.
 The drone lands, communicates with the base station, and specifies the module to swap in.
@@
 Finally, the drone flies off with the new module for its new task.
-Figure 1. RASP autonomous payload reconfiguration to execute user specified task
-Our approach of modularizing a drone’s sensors and actuators offers a number of advantages compared to a static drone with all sensors and actuators attached.
-Much like single-lens reflex (DSLR) cameras with interchangeable lenses, 1) modularization allows for the creation of an ecosystem of sensors and actuators.
-This allows consumers to purchase only the sensors and actuators they need and improve drone reusability, expendability, and sustainability.
-A static drone may not have all the sensors and actuators you may need or want, which requires you to purchase a completely new drone, rather than just the sensor or actuator.
-2) Layering and modularization enables you to evolve the drone, sensors, and actuators independently.
-New sensors and actuators purchased will still be compatible with your current drone as long as the interface remains the same and vice versa.
-3) A drone carrying a single sensor or actuator can be designed much smaller, more agile, less noisy, and more suitable for closed environments, such as indoors.
-We believe RASP to be one crucial step towards general drone platforms capable of aiding in many daily activities, much like the role of smartphones.
-To demonstrate this, we prototype and show how RASP could be integrated into a personal assistant system that also leverages static sensors in the environment (cameras) in conjunction with foundational models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting.
-For example, a drone could carry a temperature sensor to determine the warmest place to sit, before swapping in a payload module to carry a snack to an occupant.
-To the best of our knowledge, this deployment is the first end-to-end architectural, mechanical, and software design to autonomously control and reconfigure drones for versatile tasks leveraging foundational models.
+RASP is an important step towards FMs and LLMs that can dynamically interact with occupants and the environment and addresses all of the challenges mentioned previously:
+1.
+Modularization enables easy integration of new applications.
+Much like single-lens reflex (DSLR) cameras with interchangeable lenses, modularization 1) allows for the creation of an ecosystem of sensors, actuators, and applications.
+Consumers can therefore purchase only the sensors and actuators for the applications they need and improve drone reusability, expendability, and sustainability.
+A static drone may not have all the required sensors and actuators, which would require purchasing a completely new drone.
+Layering and modularization also enables 2) evolving the drone and enabling new applications independently.
+New sensors and actuators purchased for new applications will still be compatible as long as the interface remains the same.
+Drones carrying a single sensor or actuator 3) can be designed much smaller, more agile, less noisy, and more suitable for closed environments, such as indoors.
+2.
+Actuating a drone enables localized sensing and actuation.
+A reconfigurable drone platform enables FMs to dynamically specify what sensing or actuation modality at which location, without relying entirely on static sensors that may not have been deployed.
+3.
+Drones can achieve full spatial coverage while allowing for sparser deployments.
+Because FMs can actuate a drone to any location, there no longer is a need for dense deployments for all types of sensors.
+Instead, a single drone can be used to service an entire space.
+To demonstrate the utility of RASP for LLMs and FMs interfacing with physical environments, we prototype and show how RASP could be integrated into a personal assistant system that leverages static sensors deployed throughout the environment (cameras) in conjunction with foundation models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting.
+While there are several existing configurable drone platforms (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023), these works focus on the physical design and control of the drone rather than the sensing, analysis, and actuation capabilities.
+Moreover, these existing works require manual reconfiguration by humans, while RASP autonomously reconfigures the sensing and actuation capabilities on-the-fly.
+To the best of our knowledge, our work is the first to propose a drone platform with reconfigurable sensing to address challenges in coverage, localized sensing, and dense deployments for enabling general LLM and FM interactions with our physical environments.
 Our contributions are as follows.
 1.
-We create and implement RASP, a modular and reconfigurable sensing and actuation platform that allows a single drone to adapt to a wide range of scenarios.
-RASP allows drones to dynamically change their onboard sensing and actuation modules in under $25$ seconds.
+We propose and implement RASP, a modular and reconfigurable sensing and actuation platform that enables FMs to adapt the same drone to a wide range of scenarios through physical spaces.
+RASP equips FMs and LLMs with drones that can dynamically change their onboard sensing and actuation modules in under $25$ seconds.
 2.
 To realize RASP, we propose a three-layer architecture that physically replaces onboard modules, maintains data connection lines, and enables drones to communicate with sensing and actuation modules through a unified interface.
+Unlike existing reconfigurable drone platforms that focus on drone control and design, and require manual intervention to swap components, RASP focuses on systems and methods to autonomously reconfigure the drone’s onboard sensing and actuation capabilities, without human intervention.
 3.
-To show the importance of RASP for future ambient intelligent systems, we prototype and integrate RASP into a personal assistant system that takes advantage of static sensors in the environment (cameras), large language and vision language models, and the mobility of a drone to realize a wide range of useful applications in smart home, lab, office, and indoor settings.
+We demonstrate how RASP can augment LLMs and FMs and easily enable new applications throughout our environments, beyond the capabilities of common IoT smart devices, by prototyping a personal assistant system that leverages both static sensors (cameras) and the mobility and flexibility of a reconfigurable drone.
+Our deployments in realistic scenarios demonstrate that RASP can improve the success rate of a diverse array of tasks by up to $85\%$.
 2. Related Works
+1.
+Language and foundation models.
+LLMs and FMs have seen a surge in usage and research due to their powerful capabilities in allowing computers to understand and interact using natural human modes of communication in an unprecedented manner.
+While most of these works focus on generating language, text, and digital media, there is a growing trend on adapting them for autonomous systems that interact with the physical world.
+Many such works target robotic platforms (Vemprala et al., 2023), including drones and robot vacuums, that only have an egocentric view of the vicinity around them.
+Works that leverage FMs and LLMs for interacting with larger spaces, namely smart homes, generally focus on actuating common smart appliances to better respond to the needs of occupants (King et al., 2023; Nie et al., 2022b, a).
+These works, leverage LLMs to create a natural language interface between humans and their environments, often leveraging internet-connected devices and targetting applications that are already widespread (e.g., television or air conditioning control).
+Our work focuses on enabling new applications and interactions that LLMs and FMs can have with occupants and their environments scalably without being limited to the constraints imposed by the sensors and functionalities implemented and hard coded into the devices found throughout the environment.
+2.
+Reconfigurable sensing and drone platforms.
 There are many reconfigurable and modular sensing platforms in existence.
@@
 Several platforms leverage the Berkeley TinyOS operating system (Levis et al., 2005), which allow developers to develop extremely long-lasting applications with great flexibility (Culler et al., 2002; O’Flynn et al., 2005; Edmonds et al., 2005; Harte et al., 2007; Mikhaylov et al., 2013; Mikhaylov and Huttunen, 2014).
-Despite the prevalence of reconfigurable and modular sensing platforms, there currently does not exist any platform that allow drones to reconfigure and adapt its onboard sensors and actuators automatically.
-The idea of a smart home and other smart environments has existed for many decades.
-There are many smart home systems developed by research groups (Das et al., 2002; Mozer, 1998; Helal et al., 2005) and commercial products from companies like Google (LLC, 2016), Amazon (Amazon.com, 2016), and Apple (Inc., 2020).
-Most of these leverage static sensors and actuators at specific parts of the home to sense and respond to home or human activity.
-Though there are a few up and coming products (e.g., Ring Always Home Drone (rin, 2023)) and research works (Xia et al., 2021) that leverage drones for indoor and home scenarios, more research is needed in all areas (e.g., control, localization, perception, understanding) to realize a full drone-based system that can operate autonomously and take advantage of the sensing and compute from static sensors throughout indoor environments.
+There are a few reconfigurable drone platforms, but most focus on enabling flexible physical design and control of drones (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023).
+For example, (da Silva Ferreira et al., 2020) allows operators to reconfigure the number of rotors the drone with locking mechanisms that connect rotor modules.
+These works typically do not focus on sensing and actuation reconfiguration, unlike RASP.
+Additionally, they typically require a person to manually reconfigure the drone, while we introduce mechanisms that enables human-free configuration of the sensing and actuation.
 (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.
-Language and foundational models have seen a surge in usage and research due to their powerful capabilities in allowing computers to understand and interact using natural human modes of communication in an unprecedented manner.
-There are a few works that leverage language models as an interface for people to interact with their smart homes (King et al., 2023; Nie et al., 2022b, a) or drones and other robotics (Vemprala et al., 2023).
-However, to the best of our knowledge there are no works that attempt to realize a smart drone-based system for indoor settings that jointly utilizes the strengths of language models, static infrastructure, and a reconfigurable drone platform to perform a diverse range of useful tasks for occupants.
-3. Modular and Reconfigurable Sensing and Actuation Drone Platform
+3. Modular and Reconfigurable Sensing and Actuation Drone
 Figure 2. RASP architecture.
-In this section, we introduce the key components of RASP, our reconfigurable and modular sensing and actuation platform and the drone system that it is built upon.
 Figure 2 shows the architecture of RASP.
 At a high level there are three major conceptual architectural components: the ground station or mechanical layer, electrical connection layer, and the software layer.
-These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 5), sensor/actuation modules (Figures 3c-d and 4), unified module interface (UMI) (Figure 6), the carrier board (Figure 3b), and the physical drone (Figure 3a).
+These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 4), sensor/actuation modules (Figures 3c-d), unified module interface (UMI) (Figure 5), the carrier board (Figure 3b), and the physical drone (Figure 3a).
 3.1. Architecture
 Mechanical Layer.
-To enable reconfigurability, there needs to be a mechanism to physically swap out different sensor modules; the mechanical layer is responsible for this.
-The drone will land on the ground station, which will physically detach the current module onboard the drone and attach the newly requested module.
+On a functional level, the mechanical layer separates RASP from existing reconfigurable drone platforms by enabling the autonomous swapping of sensor and actuator modules.
+The drone lands on the ground station, which will physically detach the current module and attach the newly requested one.
+We propose novel swapping, including an augmented funnel shaped landing structure, to enable autonomous human-free swapping of modules as discussed in Section 3.2.
 Electrical Connection Layer.
-Each sensor and actuation module needs to maintain data, communication, and power to the drone, despite being detachable; The electrical layer maintains these connections.
-We leverage magnets to accomplish this, maintaining physical connection between the module and the drone during a mission, while still allowing the mechanical layer to easily swap out modules.
+This layer maintains the data, communication, and power connection between each sensor or actuation module with the drone, while being both detachable yet reliable mid mission.
+Existing reconfigurable drone platforms (da Silva Ferreira et al., 2020) generally incorporate slotted mechanical mechanisms.
+While these pins can maintain connection under great amounts of force, they are difficult to attach and remove, requiring a person to manually connect or disconnect.
+Attaching often requires precisely aligning slots or pins and applying acute pressure at junction points.
+Instead, we leverage magnets to maintain connection.
+Attaching a magnet-based connector only requires moving the module close to the drone’s connector before the magnetic pull aligns all connections in place.
+Removing the module can be achieved without needing to apply acute force directly at the junction.
+Both of these characteristics allow magnet-based connectors to be more conducive for autonomous reconfigurable systems than purely mechanical connectors.
+We discuss in more detail in Section 3.2.
 Software Layer.
-In addition to maintaining physical and electrical connections, the platform needs to adapt to new modules on-the-fly even if they are swapped out.
-The software layer provides standardized protocols and libraries that allow the drone to discover, configure, control, and receive data from attached modules.
+We propose standardized protocols and software libraries on top of the mechanical and electrical connection layers that allow the drone to discover, control, and receive data from attached modules on-the-fly.
 3.2. Platform Components
@@
 Drone platform.
-We designed the drone based on the open-source Crazyflie drone (cra, 2023), as shown in Figure 3.
-We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, not just in open outdoor spaces.
+We build RASP off the open-source Crazyflie drone (cra, 2023), as shown in Figure 3.
+We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, especially indoors.
 To increase its payload, we replaced its brushed motors (19000 Kv powered by a 1-cell battery) with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery).
-We also printed a protective guard, not only to protect its propellers, but to protect people and other objects from its propeller blades.
-The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie drone.
+We also printed a protective guard to protect its blades from objects and people.
+The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie.
 Carrier board.
@@
 Table 1. RASP’s supported sensors/actuators and their mass, power consumption, and cost.
-Figure 4.
-a) 3D structure attached to modules to allow the RASP to easily pick them up, b) and c) are displacement plots from stress analyses; the corresponding lowest factors of safety (FOS) are 1.180 and 4.664.
 Sensor, Actuation, and Recharging Modules.
 RASP comes with a collection of sensor and actuation modules (full list in Table 1).
-These modules are attached or detached from the drone by the ground station/mechanical layer.
+These modules are attached or detached from the drone by the ground station/mechanical layer and comes with a 3D printed structure that increases the surface for the ground station to pick up modules.
 The actuation module, shown in Figure 3d, is a container structure with a motor that opens and closes a hatch.
@@
 The recharging module is the default module that’s attached to the drone each time the drone finishes a task, returns to the ground station, and has no other pending tasks.
-To make it easier for the ground station to pick up a module, we 3D printed and attached a structure shown in Figure 4.
-We performed force analysis (using SOLIDWORKS SimulationXpress) to select the optimal structure and thickness for the frame that is as light as possible to minimize extra weight while being sturdy enough to avoid being deformed when the gripper holds it.
-We measured that the grippers applies a 5 N force onto the sides of the frame when it grips the frame and a 10 N at the bottom of the frame when it brings the module down to the conveyor belt.
-To avoid bending at these forces, our force analysis showed when the frame is 1.5mm thick, it had a factor of safety (FOS) greater than 1 in both cases, indicating that the frame was able to withstand the forces applied.
-The mass of this frame is 11.97 g, which is negligible when compared to the rest of the drone (344.7 g).
 Figure 4. Different landing platforms designed and benchmarked. The average landing time took $7.8$ seconds.
@@
 We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023) because it provides enough space for the drone to land, as well as enough mechanical control to remove onboard sensors and attach new sensors onto the drone.
-We designed and benchmarked three different grippers with different slots and grooves and measured the offset from the center of the parked drone on the ground station.
-The optimal design, shown in Figure 3e, helps the module slide into a centered position even if the module is slightly offset because of errors in actuating the conveyor to move the module into place.
+The gripper in Figure 3e removes and attaches modules onto the drone.
 The ground station contains two additional components: the communication hub and the landing station.
@@
 In future work, we plan to leverage more complex computer vision models to automatically determine the position and orientation of the drone without needing to add additional markings.
+Landing Platform.
 Removing an attached sensor and ensuring that the new sensor is securely fastened requires the drone to land precisely at the specific location where the interchange occurs.
@@
 Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023).
-As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module.
-We test and benchmark two designs for the landing station, as shown in Figure 5c.
-One design is a normal funnel (Figure 5c-top).
-The second design contains grooves to help the drone slot into place (Figure 5c-bottom); the base of the drone also has protruding grooves that match the grooves in the funnel.
+As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module (Figure 4c-top).
+To further improve alignment, we propose a new design that includes additional grooves to further improve module and drone alignment (Figure 4c-bottom).
+We benchmark both the vanilla funnel and the grooved funnel landing platform design, as shown in Figure 4d.
 For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar.
-However, the grooved design corrects for greater drone misalignments from the platform (higher maximum tolerable orientation and offset errors in Figure 5d).
-As such, we adopt the grooved design into RASP.
+However, the grooved design corrects for greater drone misalignments from the platform, as reflected in the higher maximum tolerable orientation and offset errors.
+As such, we adopt our proposed grooved funnel design into RASP.
 Figure 5. Universal module interface for connecting and disconnecting sensors and actuators, as well as recharging the drone.
@@
 The UMI has the following design features.
-Standardized 24-pin magnetic connections: Each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, and the carrier board has another set that matches.
-The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer if a module needs to be switched.
-The spring-loaded pins enables electrical connections to the sensors and actuators, while allowing the ground station to charge the drone.
-Sensor and actuator drivers: The drone needs to support a variety of different sensors and actuators, each of which may be interfaced with differently.
-To support this, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC.
+Standardized 24-pin magnetic connections: Unlike existing reconfigurable drone and sensing platforms that use mechanical connectors, each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, with corresponding connectors on the carrier board.
+The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer.
+While mechanical pins and slots are the standard for maintaining stable connections, they are difficult to attach and remove without human intervention.
+Magnets allow attaching and removing modules with much less precision, with enough force to maintain connection mid-mission.
+Magnetic connections only require placing the module in the vicinity of the drone to attach the module and diffused force to remove, compared to standard mechanical connectors.
+Sensor and actuator drivers: To support a variety of sensors and actuators that may each be interfaced differently, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC.
 We also provide developers with a driver template that allows them to easily incorporate new sensors and actuators into the RASP ecosystem.
@@
 To enable this feature, we include a non-volatile memory chip on each module (M24C02 (STMicroelectronics, 2017) 2Kb I2C bus electrically erasable programmable read-only memory (EEPROM)), which stores the identity of the module.
-The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver from a repository.
+The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver.
 4. Towards a Fully Autonomous Drone-based Personal Assistant.
 Figure 6. System architecture of intelligent assistant with RASP.
-A drone-based personal assistant is one of the core ambient intelligent systems that we see RASP enabling.
-For several decades now, the number of smart devices, sensors, and actuators deployed throughout our homes, offices, and environments has been growing exponentially (Forecast, 2021).
-These static sensing devices can provide a coarse-grained and localized picture of events and conditions within the environment; obtaining a fine-grained high resolution picture would require dense deployments of sensors, covering every square inch of the environment, which is impractical.
-We believe that adapting RASP, drones, and other mobile robotics into smart environments and ecosystems could enable fine-grained sensing and actuation at the tip of our fingers without the need for dense static deployments.
-Figure 7 shows the architecture and workflow of our prototype drone-based personal assistant.
+In this section, we detail the prototype of RASP into a personal assistant application that has information about the space through a deployment of static sensors (namely cameras).
+Sections 4.1 and 4.2 introduce methods for adapting visual language models to analyze and understand sensor data from the environment, while Sections 4.4 and 4.3 introduce how FMs can integrate RASP into their pipeline.
+Figure 6 shows the workflow of our prototype.
 Static cameras in the environment provide a high-level and coarse-grained view of the environment; in future work and systems, we envision RASP being compatible with other static sensors, beyond cameras.
@@
 The time series data from the temperature sensor will then be analyzed by another language/foundational model to make the final determination.
-There are four classes of tasks that our system can accommodate.
+There are four classes of tasks that we explore in this deployment, multiple of which are currently not commonly packaged into existing IoT smart devices:
 Object/Location Identification.
@@
 f) camera module used in ceiling camera network.
-We use images from the camera network to guide and control the drone.
-We attach an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3.
+We use images from the camera network to guide the drone by attaching an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3.
 To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest.
@@
 Moreover, as shown in Figure 10, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches.
-Using additional markers on top of the drone and the straight line path to navigate is not ideal for generalizing to other drones and environments that are more winding.
-The focus of this portion of our work is to demonstrate the utility of RASP in an important application (personal assistant) and not specifically drone detection, navigation, or control.
+While there are more practical approaches for drone navigation, the focus of this work is to demonstrate the utility of RASP to LLMs that interact with the physical world.
 Hence, we leave these aspects to future work.
 5. Deployment
-We deployed the RASP based personal assistant into an office/lab setting as shown in Figure 11e.
-We used a network of 4 cameras in a uniform grid.
-As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people walking and moving through the space.
+We deployed our RASP based personal assistant into an office/lab setting as shown in Figure 10e, using a network of 4 cameras.
+As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people moving through the space.
+The goal of this deployment is to demonstrate improvements in task completion rate RASP provides to FMs for more general and less structured applications.
 5.1. Benchmarking
@@
 This both reduces false positives and improves overall accuracy.
-Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall) and precision by incorporating RASP.
+Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall).
+In total, integrating RASP with FMs improved the task success rate by $85\%$.
 Scenario # of user prompts # of executions Execution
@@
 The previous section demonstrated how RASP could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting.
-Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy these multi-layered and more complex tasks.
+Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy multi-layered tasks.
 Common tasks that require RASP to reconfigure the drone mid-mission come in three different flavors: 1) actuate + actuate: multiple individual actuation tasks aggregated into one task (e.g., “bring me my medicine AND a snack”), 2) sense + sense: a task that involves multiple sensing modalities (e.g., “find me the coolest (temperature) place to sit out of the sunlight (light)”, 3) sense + actuate: tasks that involve performing actuation in response to sensing (e.g., “place rat poison (actuation) in dark areas (sensing)”).
@@
 Here, we placed the kettle at location three.
-Takeaways.
-We demonstrated how RASP could be used to reconfigure a drone, mid-mission, for tasks that require multiple sensors and/or actuators in a smart building setting with a decent success rate.
-However, a truly practical system still needs future improvements in drone landing (causes failed module swaps), methods for analyzing sensor data, and fine-grained control of drones.
 5.3. In-the-Wild Deployment
@@
 6. Discussion
-The focus of this work is to realize an autonomous reconfigurable and modular sensing and actuation platform drone platform that can adapt its sensing and actuation capabilities depending on the task at hand.
-As such, a single drone can be adapted for multiple tasks.
-We demonstrate the utility of such a platform for ambient intelligent systems by integrating it with a drone-based personal assistant that we envision will become commonplace in future homes, offices, and buildings.
-Specifically, we demonstrate the potential of leveraging static sensors in the environment and foundational models (e.g., visual-language models) in conjunction with a powerful and adaptable mobile platform in smart building, home, office, and indoor environments.
-To realize a fully autonomous drone-based personal assistant and other intelligent systems that can fully take advantage of RASP, there are several directions that need more exploration.
-Robust path planning, navigation, and search for indoor environments.
-As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented on the drone that were not able to capture and detect small objects of interest in a cluttered home/indoor setting.
-More work into exploration and detecting these small objects of interest is needed.
-Penetrative AI, foundational models and leveraging diverse sensors in the environment.
+The focus of this work is to demonstrate how a novel autonomous reconfigurable sensing and actuation drone platform is a critical step towards FMs, LLMs, and AGI that can interact, respond, and actuate our physical environments seamlessly, much like what has been achieved in the digital domain.
+To progress further, there are several directions that warrant further attention, which we discuss next.
+Robust drone path planning, navigation, and search for indoor environments.
+As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented that could not capture and detect small objects of interest in a cluttered home/indoor setting.
+More work into exploring, detecting, and navigating to these small objects of interest is needed.
+Penetrative AI, foundation models and leveraging diverse sensors in the environment.
 In this work, we leveraged cameras in conjunction with language to understand and interpret the environment.
-However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundational models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing our platform and case study to all types of scenarios.
-Drone noise reduction, smaller sizes, and power efficiency.
-Drones and quadcopters are very loud because of its propellers.
-Having them comfortably co-exist with people in indoor environments requires more research into the design of them to reduce their size and noise, while prolonging their battery life and enabling automatic charging while docked.
+However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundation models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing FMs and RASP to more scenarios.
 7. Conclusion
-We propose RASP, a modular and reconfigurable sensing and actuation platform that allows small drones to swap sensing and actuation capabilities dynamically.
-RASP allows a single drone to reconfigure itself to satisfy a diverse range of tasks that is commonly not possible with current drones.
-Current industrial drones are typically designed to satisfy a single application domain, and consumer drones are currently mostly used by hobbyists for entertainment.
-We demonstrate the diverse range of tasks RASP can enable by integrating RASP into a personal assistant system that also leverages static cameras to enable a wide range of useful tasks that were previously not possible.
-RASP is a step towards generalizing the use and presence of drones in daily life, much like how smartphones have become ubiquitous.
-Acknowledgements.
-This research was partially supported by COGNISENSE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA, as well as the National Science Foundation under Grant Number CNS-1943396.
-The views and conclusions contained here are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Columbia University, NSF, SRC, DARPA, or the U.S.
-Government or any of its agencies.
+We propose RASP, a modular and reconfigurable drone sensing and actuation platform that allows FMs and LLMs to dynamically to swap sensing and actuation capabilities on-the-fly.
+We demonstrate through a real deployment of a personal assistant application that RASP can improve the successful completion of complex tasks throughout our physical spaces by up to $85\%$.
+This drastic improvement stems from RASP’s ability to provide localized sensing and actuation in a large space even without large deployments of sensors and devices, as is commonly seen today.
+RASP is a critical step towards FMs and LLMs that can naturally interact and actuate the physical environment, just as they have shown in many applications in the digital domain.

### v3

@@
-RASP: A Drone-based Reconfigurable Actuation and Sensing Platform for Engaging Physical Environments with Foundation Models
+FlexiFly: Interfacing the Physical World with Foundation Models Empowered by Reconfigurable Drone Systems
 Abstract.
-Foundation models and large language models have shown immense human-like understanding and capabilities for generating text and digital media.
-However, foundation models that can freely sense, interact, and actuate the physical world like in the digital domain is far from being realized.
-This is due to a number of challenges including: 1) being constrained to the types of static devices and sensors deployed, 2) events often being localized to one part of a large space, and 3) requiring dense and deployments of devices to achieve full coverage.
-As a critical step towards enabling foundation models to successfully and freely interact with the physical environment, we propose RASP, a modular and reconfigurable sensing and actuation platform that allows drones to autonomously swap onboard sensors and actuators in only $25$ seconds, allowing a single drone to quickly adapt to a diverse range of tasks.
-We demonstrate through real smart home deployments that RASP enables FMs and LLMs to complete diverse tasks up to $85\%$ more successfully by allowing them to target specific areas with specific sensors and actuators on-the-fly.
+Foundation models (FM) have shown immense human-like capabilities for generating digital media.
+However, foundation models that can freely sense, interact, and actuate the physical domain is far from being realized.
+This is due to 1) requiring dense deployments of sensors to fully cover and analyze large spaces, while 2) events often being localized to small areas, making it difficult for FMs to pinpoint relevant areas of interest relevant to the current task.
+We propose FlexiFly, a platform that enables FMs to “zoom in” and analyze relevant areas with higher granularity to better understand the physical environment and carry out tasks.
+FlexiFly accomplishes by introducing 1) a novel image segmentation technique that aids in identifying relevant locations and 2) a modular and reconfigurable sensing and actuation drone platform that FMs can actuate to “zoom in” with relevant sensors and actuators.
+We demonstrate through real smart home deployments that FlexiFly enables FMs and LLMs to complete diverse tasks up to $85\%$ more successfully.
+FlexiFly is critical step towards FMs and LLMs that can naturally interface with the physical world.
+Embodied AI, Foundation Model Agent, Reconfigurable Drone Platform
+[…]
 1. Introduction
-Foundation models (FM) and artificial general intelligence (AGI) have the potential to bring about massive technological and social change to the physical world by discovering, performing, and continually adapting to new applications and environments without needing to manually embed or program them.
-While LLMs and FMs have been shown to have human-like understanding and digital media generation capabilities, their ability to respond to events and actuate systems in the physical world are comparatively less explored.
-There are a number of works that incorporate language models to control physical robotic systems through voice (Vemprala et al., 2023; Nie et al., 2022b), from both industry and academia, such as the Figure01 AGI robot (Figure, 2024).
-The sensing and control in such scenarios are often localized to the vicinity of the autonomous agent.
-Scaling up, there are few works that explore the use of LLMs to actuate our environments, particularly smart homes (King et al., 2023, 2024; Rivkin et al., 2023), where events and actions may occur anywhere in the space.
-These works generally focus on adapting LLMs as a human-like interface to actuate common internet-connected smart appliances (e.g., speakers, television, air conditioning, etc.).
+While there are a number of works that incorporate large language models (LLM) into robotic and egocentric systems (Vemprala et al., 2023; Nie et al., 2022b), such as the Figure01 AGI robot (Figure, 2024) and digital voice assistants, there are few works that explore the use of LLMs and foundation models (FM) to actuate our physical environments.
+Unlike egocentric systems, where sensing, processing, and control are often localized to the vicinity of the autonomous agent, executing tasks and monitoring spaces often involves sifting through heterogeneous streams of sensing data spanning large areas of the space to detect and process events localized to a tiny fraction.
+While FMs and LLMs have shown strong performance on summarization tasks, they have difficulty completing tasks that involve processing localized areas of interest without mechanisms to help them “zoom in”.
 Much like how FMs enable general human language and sensory understanding and responses, our work explores how FMs could be used to enable a diverse range of general interactions with objects and spaces that likely cannot be actuated digitally through code.
 These interactions form a large portion of our daily interactions.
-Figure 1. RASP autonomous payload reconfiguration to execute user specified task
+Figure 1. FlexiFly enables FMs to “zoom in” to areas of interest with reconfigurable drones to better interface with physical environments.
+Figure 2. System architecture of intelligent assistant with FlexiFly.
 For example, a person may want his/her home to physically bring a snack or medicine; this would require a robotic system that could physically identify and carry the payload.
@@
 A third person in a chemical lab may want the building to monitor and notify him/her about the results of a chemical reaction that s/he started before leaving; this situation could likely be accomplished with a camera with special capabilities to detect these reactions.
-While it is possible to enable all these applications, these examples highlight two main challenges that prevent FMs and LLMs for achieving the same amount of autonomy in our physical environments as we have seen in the digital domain:
+While it is possible to enable all these applications, these examples highlight three main challenges that prevent FMs and LLMs for achieving the same amount of autonomy in our physical environments as we have seen in the digital domain:
 1.
@@
 Evolving new applications and functionality in this way is also not scalable.
-Another way to state this challenge is that the generality and types of applications available are limited by the functionality of the devices and sensors present.
 2.
 Events, tasks, and actions are often localized.
-Certain events and tasks a user may require the FM to analyze the entire space, but perform an action at a target location.
-For example, placing rat poison in areas around the home that will be most effective.
-If a specific sensing modality needed to make this judgment (e.g., moisture) is not present at a specific location, then the FM cannot obtain the information it needs to carry out the task.
-This leads to the third challenge.
+Events and tasks are often localized to small areas.
+As we will show in Section 3, an FM that is analyzing data from multiple sensors covering a large space often misses events occurring in small sections of a large scene.
+While LLMs and FMs have shown great performance in analyzing general trends, we will show that they have difficulty detecting and responding to localized events in larger sensor deployments (Section 3).
 3.
@@
 While it is common for smart spaces to deploy smart devices and sensors, new applications may require dense deployments that need to be tailored to the layout of the environment, making it difficult to achieve full coverage for each new application.
-Towards enabling general task execution by FMs in physical environments, we propose RASP, an adaptive, modular, and reconfigurable actuation and sensing drone platform that allows FMs and LLMs to adapt the same drone to diverse scenarios by efficiently switching out sensors and/or actuators depending on the situation.
-Figure 1 shows an overview of RASP.
-The drone lands, communicates with the base station, and specifies the module to swap in.
-The base station secures the drone and actuates the conveyor belt and gripper to remove and replace the onboard sensor/actuator with the requested module.
-The unified module interface (UMI) ensures that the sensor/actuator module is firmly attached, connected, and can be accessed by the drone.
-Finally, the drone flies off with the new module for its new task.
-RASP is an important step towards FMs and LLMs that can dynamically interact with occupants and the environment and addresses all of the challenges mentioned previously:
+We propose FlexiFly, a platform that addresses the challenges in enabling FMs to interface with the physical world by enabling FMs, particularly visual-language models (VLMs), to “zoom in” to localized areas of interest to obtain a higher resolution, fine-grained, and better understanding of the physical environment to carry out the task at hand.
+FlexiFly accomplishes this goal with two critical design choices, as shown in Figure 1.
+First, we propose a novel image segmentation method called Aspect Ratio Constrained K-Means (ARCK-Means) segmentation to visually “zoom in” and identify potential locations pertinent to the task.
+We show how existing state-of-art segmentation methods (e.g., Segment Anything Model (SAM) (Kirillov et al., 2023)) often result in split objects that reduce object identification and localization accuracy.
+ARCK-Means segmentation ensures that full objects are processed and improves the analysis of localized areas of interest in FMs.
+Second, we propose an adaptive, modular, and reconfigurable sensing and actuation drone platform that FMs can actuate to “zoom in” physically.
+Once the FM identifies potential locations of interest, it can select the relevant sensor or actuator to equip, before actuating the drone to analyze locations of interest up close.
+For example, to answer the question: “where is the warmest place to sit?”, the FM would identify potential locations that may indicate warmth (e.g., sunlight), before sending a drone equipped with a temperature sensor to confirm.
+While there are several existing configurable drone platforms (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023), these works focus on the physical design and control of the drone rather than the sensing, analysis, and actuation capabilities.
+Moreover, these existing works require manual reconfiguration, while FlexiFly autonomously reconfigures sensing and actuation capabilities on-the-fly.
+Our motivation for exploring the use of drones to empower foundation models interfacing with the physical world stems from our vision that humans and robotic systems will coexist in the future.
+Moreover, our design choices address the challenges mentioned previously:
 1.
@@
 Instead, a single drone can be used to service an entire space.
-To demonstrate the utility of RASP for LLMs and FMs interfacing with physical environments, we prototype and show how RASP could be integrated into a personal assistant system that leverages static sensors deployed throughout the environment (cameras) in conjunction with foundation models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting.
-While there are several existing configurable drone platforms (Schiano et al., 2022; Derrouaoui et al., 2021; Perikleous et al., 2024; Hert et al., 2023), these works focus on the physical design and control of the drone rather than the sensing, analysis, and actuation capabilities.
-Moreover, these existing works require manual reconfiguration by humans, while RASP autonomously reconfigures the sensing and actuation capabilities on-the-fly.
+To demonstrate the utility of FlexiFly for LLMs and FMs interfacing with physical environments, we prototype and show how FlexiFly could be integrated into a personal assistant system that leverages static sensors deployed throughout the environment (cameras) in conjunction with foundation models and penetrative AI (Xu et al., 2024) to satisfy a wide range of useful tasks in a home, lab, or office setting.
 To the best of our knowledge, our work is the first to propose a drone platform with reconfigurable sensing to address challenges in coverage, localized sensing, and dense deployments for enabling general LLM and FM interactions with our physical environments.
@@
 1.
-We propose and implement RASP, a modular and reconfigurable sensing and actuation platform that enables FMs to adapt the same drone to a wide range of scenarios through physical spaces.
-RASP equips FMs and LLMs with drones that can dynamically change their onboard sensing and actuation modules in under $25$ seconds.
+We propose FlexiFly, a platform that enables FMs, particularly VLMs and LLMs, to better understand and interface the physical world by identifying localized areas of interest from large scenes, equipping relevant sensors/actuators, and actuating the drone to “zoom in” up close.
 2.
-To realize RASP, we propose a three-layer architecture that physically replaces onboard modules, maintains data connection lines, and enables drones to communicate with sensing and actuation modules through a unified interface.
-Unlike existing reconfigurable drone platforms that focus on drone control and design, and require manual intervention to swap components, RASP focuses on systems and methods to autonomously reconfigure the drone’s onboard sensing and actuation capabilities, without human intervention.
+To realize FlexiFly, 1) we propose a novel image segmentation method called ARCK-Means that aids FMs in identifying localized areas of interest.
+We demonstrate that ARCK-Means improves the understanding and detection of objects in large and cluttered scenes over existing segmentation techniques, namely SAM, by reducing the amount of disjoint and split objects produced by segmentation masks.
+2) We introduce a drone-based modular and reconfigurable sensing and actuation platform that enables FMs to adapt to a wide range of scenarios.
+FlexiFly actuates the drone, equipped with task-relevant sensors and actuators, to identify areas of interest to analyze up close.
 3.
-We demonstrate how RASP can augment LLMs and FMs and easily enable new applications throughout our environments, beyond the capabilities of common IoT smart devices, by prototyping a personal assistant system that leverages both static sensors (cameras) and the mobility and flexibility of a reconfigurable drone.
-Our deployments in realistic scenarios demonstrate that RASP can improve the success rate of a diverse array of tasks by up to $85\%$.
+We demonstrate how FlexiFly can augment LLMs and FMs and easily enable new applications throughout our environments, beyond the capabilities of common IoT smart devices, by prototyping a personal assistant system that leverages both static sensors (cameras) and the mobility and flexibility of a reconfigurable drone.
+Our deployments and demo (Zhao et al., 2024) in realistic scenarios demonstrate that FlexiFly can improve the success rate of a diverse array of tasks by up to $85\%$.
 2. Related Works
@@
 2.
-Reconfigurable sensing and drone platforms.
+Reconfigurable and adaptive sensing platforms.
 There are many reconfigurable and modular sensing platforms in existence.
@@
 For example, (da Silva Ferreira et al., 2020) allows operators to reconfigure the number of rotors the drone with locking mechanisms that connect rotor modules.
-These works typically do not focus on sensing and actuation reconfiguration, unlike RASP.
+These works typically do not focus on sensing and actuation reconfiguration, unlike FlexiFly.
 Additionally, they typically require a person to manually reconfigure the drone, while we introduce mechanisms that enables human-free configuration of the sensing and actuation.
 (Xia et al., 2021) recently proposed a reconfigurable drone platform, but is only a demo abstract and leaves out many details on how it autonomously swaps modules.
+Prior works have also explored adaptive sampling strategies to optimize sensor deployment and data collection, both in traditional wireless sensor networks (Lee et al., 2011) and mobile sensing platforms like underwater vehicles (Chen et al., 2012).
+Unlike these approaches that focus on sampling optimization, FlexiFly enables on-demand sensor reconfiguration guided by foundation models.
+3.
+Embodied systems.
+Traditional robotic systems often rely on hand-crafted algorithms, making adaptation to new or unforeseen situations challenging.
+Additionally, they struggle to generalize learned behaviors across different tasks, limiting their effectiveness in real-world applications (Hu et al., 2023).
+By leveraging LLMs and FMs in robotics, including drones, the generalization problem in planning (Huang et al., 2022; Lin et al., 2023) and control (Chen et al., 2024; Yang et al., 2025; Singh et al., 2022) in different tasks is partially resolved.
+However, these works still face severe limitations for solving diverse and general tasks due to insufficient real-world interactions.
+This is primarily because traditional robotic systems rely on limited and fixed sensing capabilities, often restricted to cameras; once a robotic system is built, their hardware capabilities are hardly ever upgraded.
+Many tasks require multiple sensing modalities or may need the use of different types of sensors to effectively understand, interpret, and respond to events in the physical world.
+To resolve the aforementioned limitations, we propose FlexiFly, a system that bridges the physical world with foundation models through a modular sensing platform.
+3. Preliminaries and Challenges
+In this section, we explore the limitations of current foundation models in understanding and responding to events and tasks that require interacting with physical environments.
+Such tasks a person may ask his/her home or environment could include “where is the warmest place to sit”, “where did I leave my phone”, ”bring me a snack”, and much more.
+The growth of the Internet of Things (IoT) is exponential, and the number of internet connected sensors and devices is expected to double by 2030 (Sinha, 2024).
+It is expected that all future buildings and homes will be equipped with a wide array of different sensors and devices that will improve our efficiency, automation, personalized insights, and overall quality of life.
+We envision that foundation models will be capable of understanding data collected throughout our physical environments and act as a natural interface between humans and their built environments.
+3.1. Deployment
+To test the current limitations of FMs for realizing this vision, we created the deployment shown in Figure 8e, using a network of 4 cameras.
+Since FMs for language and vision are most mature, we focus on these modalities; in the future, we believe that foundation models capable of interpreting streams of heterogeneous sensing modalities from sensors strewn throughout the environment will be possible.
+We deploy a camera network on the ceiling of our deployment area, facing directly downwards.
+We generate floor maps and stitch together frames from individual camera views, according to (Liu et al., 2022), which we input into the Large Language-and-Vision Assistant (LLaVA) (Liu et al., 2023a), a state-of-art visual-language model (VLM), along with user specified tasks and commands.
+3.2. Types of Tasks
+We identify and explore four classes of tasks in this work that are useful for creating smarter physical spaces, but are not commonly packaged into existing IoT smart devices:
+T1: Object/Location Identification (ID).
+This set of tasks requires the system to observe the environment and identify an object or location based on the user’s command (e.g., “where is my phone?”).
+In our preliminary analysis, we look at one task in this category: “where is my phone?” and place a phone at different locations in view of the cameras.
+T2: Object/Location State (State).
+This set of tasks involves learning about the state or condition of a specific object or location.
+For example, “is my food burning” would require the FM to identify food that is being cooked and if the food is producing too much smoke.
+In our preliminary analysis, we look at one task in this category: “where is the warmest place to sit?” and artificially turn up the heat in certain areas of the room by placing concealed space heaters nearby.
+T3: Surveillance (Surv).
+These tasks differ from the previous two, which are executed only once after the command is received.
+In surveillance tasks, the system needs to continuously analyze the environment until a potential event occurs (e.g., “Let me know if any chemical in my experiment spills on the table”).
+In our preliminary analysis, we look at one task in this category: “let me know if any chemical in my experiment spill on the table?” and artificially simulate this event by knocking over a glass of colored water.
+T4: Actuation (Act). Unlike previous categories, which require sensing, actuation tasks require direct physical interactions with the environment (e.g., “bring me my medicine”).
+3.3. Preliminary Analysis and Limitations
+Figure 3.
+Preliminary study: (a) task completion rate of standard FMs leveraging VLM (camera) and dense sensor networks, compared to FlexiFly; (b) Example showing that the VLM could not detect the phone in plain sight unless “zoomed in”.
+Using this basic camera network + FM setup (labeled VLM Baseline), we attempt to satisfy one command from each of the classes of actions we identified.
+We also compared against an FM analyzing data from a deployed sensor network (labeled Sensor Baseline); here, we deployed 9 sensors (temperature for the state task and alcohol sensors for the surveillance task) in a grid pattern and interpolated between sensors to obtain a map spanning the entire space.
+We ran 70 trials for each scenario and show the success rate of identifying or completing each task in Figure 3a.
+We see that the success rate is low across the board for the VLM and sensor baseline, due to three limitations.
+Limitation 1: Physical size of environments and volume of sensor information and is large, while events and objects of interest are localized to small areas.
+A FM overseeing physical environments need to process many streams of data covering a large area, such as the space of our deployment.
+However, users are generally interested in only a small portion of the environment.
+We noticed that for the ID tasks (looking for phone) and surveillance (detecting spills), our FM could often not detect these objects and events due to the event occurring in a tiny portion of the environment and the limited resolution that can be captured.
+As shown in Figure 3b, LLaVA could not identify many objects unless we manually zoom in and reduce the camera’s area of coverage.
+Limitation 2: Coverage and resolution of sensing modalities required is limited and not practically scalable.
+The VLM could not accurately identify the warmest place to sit (T2: state task) because it leverages a sensing modality where standard vision may not perform well (temperature).
+Even when we included temperature sensors (sensor baseline), the success rate is still low.
+When we simulated the warmest place to sit close to the sensor, the success rate was high, but the sensor network cannot accurately capture the dynamics of the space at areas away from the sensors, which is where success rate decreased.
+This highlights that the density of sensors the performance and understanding an FM has about the environment.
+Additionally, while detecting the warmest place to sit may require a temperature sensor, different applications require different sensor deployments (e.g., detecting falls could use audio or vibration).
+For instance if audio is used to detect falls and the bedroom does not have a smart speaker or microphone, then it would be difficult to enable this service in the bedroom.
+Limitation 3: Actuation is restricted. An actuation task such as bringing the user a snack or medicine cannot be accomplished with only an FM along and static devices throughout the environments.
+3.4. Design Philosophy
+To address the limitations discussed previously, our design philosophy involves creating mechanisms that enable FMs to “zoom in” and sense targeted areas of interest with higher resolution.
+We tackle these limitations on two fronts.
+First, we propose a novel image segmentation technique called Aspect Ratio Contrained K-Means (ARCK-Means) to aid in identifying potential locations of interest to “zoom in”.
+Second, we propose FlexiFly, a drone platform, with modular and reconfigurable sensing and actuation, that allows FMs to pick and choose sensors and actuators depending on the task at hand.
+The FM first leverages ARCK-Means to identify areas of interest based on the current state of the environment and the input command (Section 4), before actuating the drone with the corresponding sensors to “zoom in”(Section 5).
+For example, an FM looking to answer “where is the warmest place to sit” previously could only rely on analyzing and sifting through a large amount of sensor data from static sensors throughout the environment.
+With FlexiFly, an FM identifies local areas of interest with ARCK-Means, before actuating a temperature sensor equipped drone to areas of interest to measure and compare temperature readings.
+ARCK-Means alleviates the first limitation by aiding the FM in identifying local areas of interest in large cluttered scenes.
+Our reconfigurable sensing and actuation drone platform alleviates limitation 2 and 3.
+Instead of requiring a dense deployment of static sensors (limitation 2), FMs can reconfigure and actuate the drone to the precise location it needs to sense, which reduces deployment overhead.
+Moreover, a mobile drone platform enables actuation of items in the physical environment that are more restrictive for static sensors and devices (limitation 3).
+4. Identifying Local Areas of Interest via Segmentation.
+Figure 4.
+Segmentation and clustering to break down scenes into smaller more manageable pieces for LLaVA and DINO.
+(a) Object masks after applying Segment Anything Model (SAM); (b) Extracted frames after clustering object masks based on K-Means; (c) Hierarchical clustering, and (d) ARCK-Means.
+For ARCK-Means, we constrain the aspect ratio of extracted frames to be between 0.67 and 1.5.
+Figure 5. Different clustering methods and thresholding evaluated for segmentation. Prompt 1: Describe the image in detail. Prompt 2: Is there a {object name} in the image?
+As mentioned in Section 3, when a user command arrives, we input the command along with frames stitched together from the camera network in LLaVA.
+We found that LLaVA was not able to reliably detect objects of interest in a scene when we passed in an image or a large space (e.g., a single chair in a large room); the larger the scene, the less detailed and more high-level the responses became.
+One way to alleviate this is to segment and analyze smaller portions of the full image.
+Instead of directly passing the entire scene into LLaVA, we use the state-of-art image segmentation model, Segment Anything Model (SAM), to extract key objects and smaller areas of the scene (Kirillov et al., 2023).
+Figure 4a shows an example of the object masks that are output by SAM.
+These object masks are then input into LLaVA, along with the user command, and outputs (in text) the names of potential objects in the scene that might be of interest.
+These text outputs from LLaVA and the segmented images are input into Grounding DINO (Liu et al., 2023b), a state-of-art language model for zero-shot object detection and localization, which then outputs localized bounding boxes of potential objects of interest.
+In implementing this segmentation pipeline for identifying locations of interest, we noticed a number of limitations that impacted performance, as discussed next.
+4.1. Limitations of Current Segmentation Methods
+1.
+SAM often generates segmentation masks that split objects between two different masks.
+As such, LLaVA and Grounding DINO often identify the same object multiple times at slightly different locations when directly using the masks from SAM, adding to the processing time.
+2.
+The segmented masks are not rectangular, which reduces object detection performance.
+We observed that directly inputting the segmented masks into LLaVA and Grounding DINO saw a large performance degradation.
+This performance degradation persisted even if we zero pad the segmented images into a rectangular shape.
+We suspect this is because LLaVA and Grounding DINO are typically trained with rectangular images, with standard aspect ratios (e.g., 640 x 480), in natural settings; whereas, the inputs we were trying are non-rectangular with most of the background removed (zero padded).
+4.2. Aspect Ratio Constrained K-Means (ARCK-Means) Segmentation
+To address the limitations of current segmentation methods in aiding FMs identify potential locations of interest, we propose Aspect Ratio Constrained K-Means (ARCK-Means) Segmentation.
+ARCK-Means operates on the segmented objects generated by SAM.
+To alleviate the challenge of splitting objects between different masks (limitation 1), ARCK-Means first clusters the centroids of the masks generated into $k$ clusters.
+We leverage hierarchical clustering (Murtagh and Contreras, 2012), but also benchmark against K-means (Ahmed et al., 2020).
+To ensure that whole rectangular images are analyzed by LLaVA and Grounding DINO (limitation 2), we find the minimum-area rectangle that encompasses all masks of each cluster.
+Moreover, we constrain the aspect ratio of clustered masks between to ensure aspect ratios of typical images used during training, resulting in the full ARCK-Means method.
+To accomplish this, we check if adding a new segment into the existing cluster causes the aspect ratio to fall below or above our constraints; if these constraints are broken, then we do not make the assignment.
+The resulting image segments are then analyzed by LLaVA and DINO.
+4.3. Analysis and Benchmarking
+Figure 4b-d shows an example of segmentation and clustering.
+We see that for ARCK-Means the clustered masks generated tend to be more square.
+Additionally, the cluttered table in the bottom right hand corner is fully encompassed by ARCK-Means, but both K-Means and hierarchical clustering do not capture all of the items on the table in a single segment.
+These improvements of ARCK-Means over the other clustering methods yields higher performance in detecting objects of interest, as shown in Figure 5.
+In this case and for the rest of the paper, a successful “recall” means that the Grounding DINO model was able to detect and localize the object interest in any one of the segmented clusters it was given, regardless of any additional detections.
+First, we tested two prompts with the LLaVA model and vary the size of the object with respect to the frame of one camera (Figure 5a).
+The first prompt is more general, asking LLaVA to describe the objects present.
+The second prompt is more specific, asking if the specific object of interest is present in the scene.
+We see that the more specific a prompt is, the higher the recall, which is the reason why we used more specific prompting.
+Second, we see that as the size of the object gets smaller with respect to the image frame, the recall gets smaller, since the signal-to-noise ratio of the object gets smaller.
+In Figure 5b, we compare the recall of each clustering method after clustering each scene into 5 segments, and see that ARCK-Means has the highest recall due to improvements in the masks it clusters (addressing limitation 1) and the shape of the resulting segments (addressing limitation 2); as such, we adopt ARCK-Means into the final system.
+Figure 5c shows the run time and recall of ARCK-Means as a function of the number of clusters or segments we split the scene into.
+We see that the recall levels out at around $90\%$ after $5$ clusters, while taking around $100$ms to run the full pipeline.
+Adding in more clusters does not yield significant improvements in detection, but significantly increases run time; as such, we segment the scene into five segments in the final system.
+To generate these plots, we used 47 images of indoor home, office, and lab environments.
+We took 35 images from the ADE20K scene parsing dataset (Zhou et al., 2017) as well as 12 images from our own deployment.
+We implement and run the full visual-language model pipeline (SAM, clustering, LLaVA, and DINO) on an Nvidia RTX 3090 GPU server.
 5. Modular and Reconfigurable Sensing and Actuation Drone
-Figure 2. RASP architecture.
-Figure 2 shows the architecture of RASP.
-At a high level there are three major conceptual architectural components: the ground station or mechanical layer, electrical connection layer, and the software layer.
-These conceptual architectural components are implemented across five physical platform components: the ground station (Figures 3e and 4), sensor/actuation modules (Figures 3c-d), unified module interface (UMI) (Figure 5), the carrier board (Figure 3b), and the physical drone (Figure 3a).
-3.1. Architecture
-Mechanical Layer.
-On a functional level, the mechanical layer separates RASP from existing reconfigurable drone platforms by enabling the autonomous swapping of sensor and actuator modules.
-The drone lands on the ground station, which will physically detach the current module and attach the newly requested one.
-We propose novel swapping, including an augmented funnel shaped landing structure, to enable autonomous human-free swapping of modules as discussed in Section 3.2.
-Electrical Connection Layer.
-This layer maintains the data, communication, and power connection between each sensor or actuation module with the drone, while being both detachable yet reliable mid mission.
-Existing reconfigurable drone platforms (da Silva Ferreira et al., 2020) generally incorporate slotted mechanical mechanisms.
-While these pins can maintain connection under great amounts of force, they are difficult to attach and remove, requiring a person to manually connect or disconnect.
-Attaching often requires precisely aligning slots or pins and applying acute pressure at junction points.
-Instead, we leverage magnets to maintain connection.
-Attaching a magnet-based connector only requires moving the module close to the drone’s connector before the magnetic pull aligns all connections in place.
-Removing the module can be achieved without needing to apply acute force directly at the junction.
-Both of these characteristics allow magnet-based connectors to be more conducive for autonomous reconfigurable systems than purely mechanical connectors.
-We discuss in more detail in Section 3.2.
-Software Layer.
-We propose standardized protocols and software libraries on top of the mechanical and electrical connection layers that allow the drone to discover, control, and receive data from attached modules on-the-fly.
-3.2. Platform Components
+In this section, we introduce the reconfigurable and modular sensing and actuation platform for drones, which FMs can actuate to identified locations of interest (Section 4).
 Figure 6. a) Physical drone system, b) Carrier board, c) Several sensor modules, d) Actuation module, e) Gripper for swapping sensor and actuation modules
+5.1. Challenges and Design
+Figure 7. Different landing platforms designed and benchmarked. The average landing time took $7.8$ seconds.
+The primary challenges in realizing an automated and reconfigurable sensing platform for drones is two fold.
+1.
+Drone and module connector.
+To allow reconfiguration, the sensor and actuation modules need to be detachable from the drone’s main body.
+While there are drone platforms that can be reconfigured with different frames, wings, or motors (Derrouaoui et al., 2022; Schiano et al., 2022; da Silva Ferreira et al., 2020), the vast majority leverage mechanical connectors with locking mechanisms that often requires additional force and complex mechanisms to fasten and remove (e.g., grippers with joints much like human hands).
+Instead, we leverage a fully magnetic connector.
+The advantage of this design choice is two-fold.
+First, it simplifies the design of the mechanisms for removing and fastening new modules on the mechanical layer.
+Second, less force is required to fasten and remove the module.
+Compared to a mechanical design, such as (da Silva Ferreira et al., 2020), that requires the application of force with several mechanisms at multiple locations, potentially damaging the module, connector, or drone, a magnetic connector enables gripper to bring a new module within vicinity of the drone before the magnet automatically aligns the pins and fastens the module in place.
+Removing the module follows a similar procedure.
+We tested the success rate in swapping a module on a drone with both magnetic (Amazon, 2024) and standard PCB mezzanine connectors (Amphenol ICC, 2020a, b) using our gripper (Figure 6e) that actuates vertically to bring modules up and down.
+We ran 10 trials and found that our system could swap in new modules successfully every time when the connectors are magnetic.
+However, our system failed at swapping modules with the standard mezzanine connectors because our mechanical grippers could not apply enough force to remove nor attach modules.
+2.
+Alignment.
+Swapping in new modules onto the drone can only occur successfully if the drone is oriented correctly with the gripper after landing.
+This challenge is not directly related to the well studied problem of drone landing (Wang et al., 2022), which aims to guide the drone to the landing platform before descending.
+As the drone descends closer to the ground, nonlinearities in drone stability (ground effect) (Matus-Vargas et al., 2021), for which there is no promising method to compensate, can cause the drone to become misaligned.
+Instead, we focus on the design of the landing platform that automatically calibrates the orientation of the drone as it touches down.
+Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023).
+As long as the drone lands within the opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module (Figure 7c-top).
+To further improve alignment, we propose a new design that includes additional grooves to further improve module and drone alignment (Figure 7c-bottom).
+We benchmark both the vanilla funnel and the grooved funnel landing platform design, as shown in Figure 7d.
+For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar.
+However, the grooved design corrects for greater drone misalignments from the platform, as reflected in the higher maximum tolerable orientation and offset errors.
+As such, we adopt our proposed grooved funnel design into FlexiFly.
+5.2. Platform Components and Implementation Details
+This section discusses the implementation of the rest of the reconfigurable platform.
 Drone platform.
-We build RASP off the open-source Crazyflie drone (cra, 2023), as shown in Figure 3.
-We chose the lightweight (27g) Crazyflie as our foundation due to its compact size, making it easier to operate in all environments, especially indoors.
-To increase its payload, we replaced its brushed motors (19000 Kv powered by a 1-cell battery) with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery).
-We also printed a protective guard to protect its blades from objects and people.
-The drone can hover for nearly 4 minutes, which is comparable to the Crazyflie.
+We build FlexiFly off the open-source Crazyflie drone (cra, 2023), as shown in Figure 6.
+We replaced its brushed motors with more powerful brushless motors (3800 Kv powered by a 4-cell 850mAh battery) .
+To improve payload capacity and ensure operational safety in areas with human presence, we designed and fabricated a enclosed drone frame using lightweight foaming polylactic acid (LW-PLA), as shown in Figure 6a.
 Carrier board.
-The carrier board, which is attached to the base of the drone, provides the physical data, power, and communication connections between the drone platform and each sensor/actuation module.
-We implement the carrier board on a lightweight \$15 Raspberry Pi Zero 2W (rpi, 2023), which has connections for one sensor or actuation module.
+The carrier board, attached to the base of the drone, provides the physical data, power, and communication connections between the drone platform and each sensor/actuation module.
+We implement the carrier board on a lightweight \$15 Raspberry Pi Zero 2W (rpi, 2023), which has magnetic connections for one sensor or actuation module.
 Total Module Fly Drone Module Cost
@@
 Actuator 394.8g 50.1g 3m08s 235.1W 1.22W \$7
-Table 1. RASP’s supported sensors/actuators and their mass, power consumption, and cost.
-Sensor, Actuation, and Recharging Modules.
-RASP comes with a collection of sensor and actuation modules (full list in Table 1).
-These modules are attached or detached from the drone by the ground station/mechanical layer and comes with a 3D printed structure that increases the surface for the ground station to pick up modules.
+Table 1. FlexiFly’s supported sensors/actuators and their mass, power consumption, and cost.
+Sensor and Actuation Modules.
+FlexiFly comes with a collection of sensor and actuation modules (full list in Table 1).
+These modules are attached or detached from the drone by the ground station and comes with a 3D printed structure that increases the surface for the ground station to pick up modules.
 The actuation module, shown in Figure 6d, is a container structure with a motor that opens and closes a hatch.
 Small items (e.g., medication, candy, pet food, etc.) can be loaded into this module for the drone to deliver.
-Additionally, we create a module that recharges the battery of the drone, which gets attached and removed just like other sensors and actuators.
-Unlike other standalone sensors and actuators, the recharging module is connected to the wall-powered ground station and charges the drone’s battery at a rate of 1.7 A.
-At this rate, the fully depleted drone can be fully charged in 30 minutes.
-The recharging module is the default module that’s attached to the drone each time the drone finishes a task, returns to the ground station, and has no other pending tasks.
-Figure 4. Different landing platforms designed and benchmarked. The average landing time took $7.8$ seconds.
 Ground station and automated takeoff and landing.
-The ground station primarily spans the mechanical layer, consisting of a gripper mechanism and a conveyor belt to position modules to swap in/out.
-We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023) because it provides enough space for the drone to land, as well as enough mechanical control to remove onboard sensors and attach new sensors onto the drone.
+The ground station consists of a gripper mechanism and a conveyor belt to position modules below the drone.
+We created the platform leveraging the chassis of the open-source Ender-3 3D printer (end, 2023).
 The gripper in Figure 6e removes and attaches modules onto the drone.
-The ground station contains two additional components: the communication hub and the landing station.
-The communication hub is the central processor that controls the swapping process.
-Drones communicate with this hub prior to landing to specify which module it requires and to initiate the swapping process.
-The drone then lands at the landing station before the mechanical layer begins the swapping process.
-The communication hub guides the drone into position for landing using two cameras facing up on top of the ground station (Figure 4b).
+To land, the ground station guides the drone using two cameras facing up on top (Figure 7b).
 We also print and attach two ArUco markers to the bottom side of the drone to easily detect the position of the drone and its orientation.
@@
 In future work, we plan to leverage more complex computer vision models to automatically determine the position and orientation of the drone without needing to add additional markings.
-Landing Platform.
-Removing an attached sensor and ensuring that the new sensor is securely fastened requires the drone to land precisely at the specific location where the interchange occurs.
-However, precise drone landing and alignment is extremely difficult because nonlinearities in drone stability become more pronounced as the drone moves closer to the ground (ground effect) (Matus-Vargas et al., 2021).
-Rather than creating a flat landing pad that is similar to the landing pads for helicopters, we take inspiration from the Ring Always Home Drone, which uses a funnel (rin, 2023).
-As long as the drone lands within the larger radius of opening of the funnel, the drone will automatically slide towards the bottom of the funnel with an opening that latches onto the onboard module (Figure 4c-top).
-To further improve alignment, we propose a new design that includes additional grooves to further improve module and drone alignment (Figure 4c-bottom).
-We benchmark both the vanilla funnel and the grooved funnel landing platform design, as shown in Figure 4d.
-For each design, we had the drone land and swap modules 75 times and found that the swap rates are fairly similar.
-However, the grooved design corrects for greater drone misalignments from the platform, as reflected in the higher maximum tolerable orientation and offset errors.
-As such, we adopt our proposed grooved funnel design into RASP.
-Figure 5. Universal module interface for connecting and disconnecting sensors and actuators, as well as recharging the drone.
-Unified Module Interface (UMI).
-The UMI (Figure 5) spans both the electrical connection layer and the software layer, giving the carrier board on the drone a standardized way to communicate with a variety of sensing and actuation modules.
-The UMI has the following design features.
-Standardized 24-pin magnetic connections: Unlike existing reconfigurable drone and sensing platforms that use mechanical connectors, each sensor and actuation module in RASP comes equipped with two magnetic connectors each with 12 spring-loaded pins, with corresponding connectors on the carrier board.
-The magnets allow each module to remain attached and connected to the drone during flight, while being pliable enough to be easily removed by the mechanical layer.
-While mechanical pins and slots are the standard for maintaining stable connections, they are difficult to attach and remove without human intervention.
-Magnets allow attaching and removing modules with much less precision, with enough force to maintain connection mid-mission.
-Magnetic connections only require placing the module in the vicinity of the drone to attach the module and diffused force to remove, compared to standard mechanical connectors.
-Sensor and actuator drivers: To support a variety of sensors and actuators that may each be interfaced differently, RASP implements a library of standard drivers, as part of the software layer, that support different communication and sampling schemes, namely: I2C, SPI, UART, GPIO, and ADC.
-We also provide developers with a driver template that allows them to easily incorporate new sensors and actuators into the RASP ecosystem.
-Plug-and-play controller: Once a new module attaches to the carrier board, the drone needs to be able to start collecting data or using the actuator automatically, without needing to manually run any command.
-To enable this feature, we include a non-volatile memory chip on each module (M24C02 (STMicroelectronics, 2017) 2Kb I2C bus electrically erasable programmable read-only memory (EEPROM)), which stores the identity of the module.
-The carrier board will pull data from here every time a new module is attached and automatically downloads the corresponding driver.
-4. Towards a Fully Autonomous Drone-based Personal Assistant.
-Figure 6. System architecture of intelligent assistant with RASP.
-In this section, we detail the prototype of RASP into a personal assistant application that has information about the space through a deployment of static sensors (namely cameras).
-Sections 4.1 and 4.2 introduce methods for adapting visual language models to analyze and understand sensor data from the environment, while Sections 4.4 and 4.3 introduce how FMs can integrate RASP into their pipeline.
+6. Implementation of a Foundation Model and Drone-based Assistant
+We show how FMs can leverage our novel methods for identifying local areas of interest (Section 4) and reconfigurable drone platforms (Section 5) to better understand and interact with physical environments, with the implementation and evaluation of a personal assistant system.
 Figure 2 shows the workflow of our prototype.
-Static cameras in the environment provide a high-level and coarse-grained view of the environment; in future work and systems, we envision RASP being compatible with other static sensors, beyond cameras.
-When a user gives a voice command, the system leverages a visual-language model to analyze the scene, interpret the command, and output a list of potential locations of interest, along with the relevant sensor or actuator.
-Then, the system actuates the RASP drone with the sensor/actuator to each of the locations to complete the task.
-For example, one task or question a person may want to know is “where is the warmest place to sit”?
-A standard RGB camera would not be able to accurately make this determination without additional sensors (e.g., temperature).
-The voice command and images are analyzed by a visual-language model, which outputs potential locations where a person could sit (e.g., chairs and couches).
-Then, the system outfits the RASP drone with a temperature module and actuates the drone to each of these locations to measure temperature.
-The time series data from the temperature sensor will then be analyzed by another language/foundational model to make the final determination.
-There are four classes of tasks that we explore in this deployment, multiple of which are currently not commonly packaged into existing IoT smart devices:
-Object/Location Identification.
-This set of tasks requires the system to observe the environment and identify an object or location based on the user’s command.
-For example, the command “where is my phone” requires the visual-language model to identify potential locations for where the phone could be, before sending a drone with a camera to take a closer look and verify.
-Object/Location State.
-This set of tasks involves learning about the state or condition of a specific object or location.
-For example, “is my food burning” would require the system to identify food that is being cooked, before sending a drone with a CO2 sensor to determine the state of the food.
-Surveillance.
-These tasks differ from the previous two, as it requires the visual-language model to continuously analyze the scene until a potential event occurs, before sending the drone (e.g., “Let me know if any chemical spills on the table”).
-Actuation.
-Unlike the previous three categories, all of which require the drone to sense the environment, actuation tasks all leverage the actuation module to bring an item (e.g., medicine, candy, food) to a specific location.
-4.1. Setup and Networked Camera System
-To effectively use a camera network to observe the environment, there generally is a lengthy setup process where you need to specify the location of the cameras and have the floor map of the building to stitch together different views.
-We adopt this self localizing camera network (Liu et al., 2022) that allows users to semi-randomly place cameras on the ceiling, facing downwards perpendicular to the floor; by leveraging motion and people walking throughout daily life, the camera network automatically self-localizes each camera within the network and generates the walkable floormap of the areas observed by the cameras.
-In this way, users only need to install cameras and do not need to define a floormap or specify and calibrate based on location.
-More details of our deployment are in Section 5.
-4.2. Visual-Language Models for Interpreting Voice Commands
-Figure 7.
-Summary of prompts used to engineer the LLaVA visual-language model (left) to interpret and identify locations of interest given a voice command from a user, as well as for ChatGPT (right) to analyze time series data from the RASP onboard drone sensor.
-Figure 8.
-Segmentation and clustering to break down scenes into smaller more manageable pieces for LLaVA and DINO.
-(a) Object masks after applying Segment Anything Model (SAM).
-Extracted frames after clustering object masks based on K-Means (b), hierarchical clustering (c), and ARCK-Means (d).
-For ARCK-Means, we constrain the aspect ratio of extracted frames to be between 0.67 and 1.5.
-Figure 9. Different clustering methods and thresholding evaluated for segmentation. Prompt 1: Describe the image in detail. Prompt 2: Is there a {object name} in the image?
-Identifying Objects of Interest.
-When a user issues a voice command, we pass this command along with the full image of the environment stitched together from all cameras into the open-source Large Language-and-Vision Assistant (LLaVA) visual-language model (Liu et al., 2023a).
-For the surveillance category, we pass in two consecutive frames or segments to detect changing events.
-LLaVA outputs (in text) the names of potential objects in the scene that might of interest, as well as the suitable sensor/actuation module that the drone should use.
-Figure 7-left shows the series of engineered and task-specific prompts we leverage, which was designed based off of the highly active Brex LLM prompt engineering guide (Brex Inc., 2023).
-The LLaVA text outputs are input as a prompt into Grounding DINO (Liu et al., 2023b), a state-of-art language model for zero-shot object detection, which then outputs bounding boxes and locations of the specified objects.
-Object Segmentation and Clustering.
-However, we found that LLaVA was not able to reliably detect objects of interest in a scene when we passed in an image or a large space (e.g., a single chair in a large room); the larger the scene, the less detailed and more high-level the responses became.
-To capture small details within the scene that might be relevant to a voice command, we used the Segment Anything Model (SAM) to extract key objects and smaller areas of the scene (Kirillov et al., 2023).
-Figure 8a shows an example of the object masks that are output by SAM.
-We tried inputting these masked objects into LLaVA and then DINO one-by-one, but there were two issues we noticed.
-First, many of the masks generated often splits objects between two different masks; as such, DINO would often detect the same object multiple times.
-Second, the masks are not shaped like a regular image (e.g., a square or rectangle); we saw a large degradation in detection performance if we zero-padded the areas around the masked object into a rectangular shape.
-To rectify this, we cluster multiple segments together to ensure potentially important objects are all contained with a single rectangular segment.
-We benchmark the following clustering methods: K-means (Ahmed et al., 2020), hierarchical clustering (Murtagh and Contreras, 2012), and Aspect Ratio Constrained K-Means (ARCK-Means).
-For K-means and hierarchical clustering, we cluster the centroids of each mask from SAM into $k$ different clusters.
-Next, we find the minimum-area rectangle that encompasses all masks of each cluster.
-These $k$ minimum-area rectangles are then analyzed by LLaVA and DINO.
-One of the challenges we found with DINO is that it was not able to identify objects of interest very well if input images are very long or very wide.
-As such, we create and benchmarked a new clustering method called Aspect Ratio Constrained K-Means (ARCK-Means), which constrains the aspect ratio of the clustered masked between two values $x<y$ to maintain a more square shape.
-To accomplish this, we check if adding a new segment into the existing cluster (during hierarchical clustering) causes the aspect ratio to fall below or above our constraints (e.g., $<x$ or $>y$); If these constraints are broken, then we do not make the assignment.
-Figure 8b-d shows an example of segmentation and clustering.
-We see that for ARCK-Means the clustered masks generated tend to be more square.
-Additionally, the cluttered table in the bottom right hand corner is fully encompassed by ARCK-Means, but both K-Means and hierarchical clustering do not capture all of the items on the table in a single segment.
-These improvements of ARCK-Means over the other clustering methods yields higher performance in detecting objects of interest, as shown in Figure 9.
-In this case and for the rest of the paper, a successful “recall” means that the Grounding DINO model was able to detect and localize the object interest in any one of the segmented clusters it was given, regardless of any additional detections.
-First, we tested two prompts into the LLaVA model and vary the size of the object with respect to the frame of one camera (Figure 9a).
-The first prompt is more general, asking LLaVA to describe the objects present.
-The second prompt is more specific, asking if the specific object of interest is present in the scene.
-We see that the more specific a prompt is, the higher the recall, which is the reason why we used more specific prompting (Figure 7).
-Second, we see that as the size of the object gets smaller with respect to the image frame, the recall gets smaller, since the signal-to-noise ratio of the object gets smaller.
-In Figure 9b, we compare the recall of each clustering method after clustering each scene into 5 segments, and see that ARCK-Means has the highest recall due to improvements in the masks it clusters and the shape of the resulting segments; as such, we adopt ARCK-Means into the final system.
-Figure 9c shows the run time and recall of ARCK-Means as a function of the number of clusters or segments we split the scene into.
-We see that the recall levels out at around $90\%$ after $5$ clusters, while taking around $100$ms to run the full pipeline.
-Adding in more clusters does not yield significant improvements in detection, but significantly increases run time; as such, we segment the scene into five segments in the final system.
-To generate these plots, we used 47 images of indoor home, office, and lab environments.
-We took 35 images from the ADE20K scene parsing dataset (Zhou et al., 2017) as well as 12 images from our own deployment.
-We implement and run the full visual-language model pipeline (SAM, clustering, LLaVA, and DINO) on an Nvidia RTX 3090 GPU server.
+Static cameras in the environment provide a high-level and coarse-grained view of the environment, just like in our preliminary deployment (Section 3.1).
+When a user gives a voice command, the system leverages LLaVA and ARCK-Means segmentation (Section 4) to identify local areas of interest and the relevant sensor or actuator.
+Then, the system actuates the FlexiFly drone with the sensor/actuator to each location to complete the task.
+We carry out and evaluate this system on the four classes of tasks we identify in Section 3.2.
+Implementation Details.
+A local server runs the Ollama framework, hosts the LLMs (Llama-3.1-8B), VLM (LLaVA 1.6-8b), and Grounding DINO open-set object detection model.
+Upon user request, a snapshot is taken from the ceiling camera network and stitched into a single image.
+The server identifies required sensing modules and key areas of interest, manages pipeline execution, and sends commands to the drone.
+A validation LLM (Llama-3.1-8B) is used that ensures properly formatted outputs throughout the process.
+For sensing tasks (ID or State tasks from Section 3.2), the drone flies to each location while reading sensor values from the attached module.
+If the sensing modality involves camera-based object identification (e.g., an object or location ID task), the captured drone image is processed on our local server.
+However, for time-series data analysis, the system uses GPT-4 (OpenAI, 2023), as we found that locally hosted LLMs often produce unreliable results when generating code to analyze the sensor data.
+In an actuation task, the system attaches the actuation module with the relevant payload (e.g., snack or medicine) before flying to the destination for dropoff.
 Prompting users for more context.
@@
 If the visual-language model detects multiple potential locations, it will ask the user to clarify the location, either through voice or a web application that we implemented.
-4.3. Drone Actuation and Analysis
-Once the visual-language models determine important locations, the system attaches the relevant sensor or actuator (from LLaVA) to the RASP drone before moving the drone to each of these locations.
-In the case of an actuation task, the drone flies to the specified location, opens the hatch to drop the item of interest, before flying back to the ground station.
-In the case of a sensing task, the drone flies to each location while reading sensor values from the attached module.
-After the drone lands back at the ground station, the time series data is input into the GPT-4 language model (OpenAI, 2023) to analyze and make the final determination.
-Figure 7-right shows the prompts we used for GPT-4.
-If the sensing modality is to use a camera to identify an object (e.g., a objection/location ID task); the image from the drone is input back into the Grounding DINO model to make the final determination.
 6.1. Drone Navigation
@@
 f) camera module used in ceiling camera network.
-We use images from the camera network to guide the drone by attaching an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3.
+During flight, we use images from the camera network to guide the drone by attaching an ArUco marker to the top of the drone, just like the patterns used for landing the drone in Section 5.
 To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest.
@@
 Moreover, as shown in Figure 8, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches.
-While there are more practical approaches for drone navigation, the focus of this work is to demonstrate the utility of RASP to LLMs that interact with the physical world.
-Hence, we leave these aspects to future work.
-5. Deployment
-We deployed our RASP based personal assistant into an office/lab setting as shown in Figure 10e, using a network of 4 cameras.
-As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people moving through the space.
-The goal of this deployment is to demonstrate improvements in task completion rate RASP provides to FMs for more general and less structured applications.
+While there are more practical approaches for drone navigation, the focus of this work is to demonstrate the utility of FlexiFly to LLMs that interact with the physical world.
+Hence, we leave these aspects for future work.
+7. Deployment and Evaluation
+We deployed our system into an office/lab setting as shown in Figure 8e, just as in our preliminary study.
+The goal of this deployment is to demonstrate improvements in task completion rate FlexiFly provides to FMs for more general and less structured applications.
 7.1. Benchmarking
 Figure 9. Summary of object and event detection by the LLaVA + DINO visual-language pipeline averaged (a) and broken down by category of sensing tasks (b-d).
-Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
 Figure 10.
-Breakdown of accuracy and false positive detections by sensing task category with and without RASP.
-We see that leveraging RASP in conjunction with static cameras greatly reduces false detections and improves accuracy because the drone can get a closeup view or sense an important part of the environment that a camera alone cannot (e.g., humidity).
-Camera Only Camera + RASP
-Scenario Precision Recall F-1 Accuracy Sensor Used Precision Recall F-1 Accuracy
+Breakdown of accuracy and false positive detections by sensing task category with and without FlexiFly.
+We see that leveraging FlexiFly in conjunction with static cameras greatly reduces false detections and improves accuracy because the drone can get a closeup view or sense an important part of the environment that a camera alone cannot (e.g., humidity).
+Camera Baseline Camera + ARCK-Means only (Section 4) Camera + FlexiFly (Sections 4 and 5)
+Scenario Precision Recall F-1 Accuracy Precision Recall F-1 Accuracy Sensor Used Precision Recall F-1 Accuracy
 Object / Location Identification
-Find Phone 68.85% 84.00% 75.68% 73.00% Drone Cam 100.00% 84.00% 91.30% 92.00%
-Find Key 78.05% 80.00% 79.01% 71.67% Drone Cam 100.00% 80.00% 88.89% 86.67%
-Sit - Temperature 23.47% 92.00% 37.40% 25.24% Temperature 76.67% 92.00% 83.64% 91.26%
-Sit - Humidity 25.81% 88.89% 40.00% 28.00% Humidity 82.76% 88.89% 85.71% 92.00%
-Sit - Light 20.62% 83.33% 33.06% 22.12% Light Sensor 95.24% 83.33% 88.89% 95.19%
-Average (ID) 43.36% 85.64% 53.03% 44.01% 90.93% 85.64% 87.69% 91.42%
+Find Phone 33.33% 26.00% 29.21% 37.00% 68.85% 84.00% 75.68% 73.00% Drone Cam 100.00% 84.00% 91.30% 92.00%
+Find Key 26.67% 10.00% 14.55% 21.67% 78.05% 80.00% 79.01% 71.67% Drone Cam 100.00% 80.00% 88.89% 86.67%
+Sit - Temperature 15.91% 56.00% 24.78% 17.48% 23.47% 92.00% 37.40% 25.24% Temperature 76.67% 92.00% 83.64% 91.26%
+Sit - Humidity 20.45% 66.67% 31.30% 21.00% 25.81% 88.89% 40.00% 28.00% Humidity 82.76% 88.89% 85.71% 92.00%
+Sit - Light 20.88% 79.17% 33.04% 25.96% 20.62% 83.33% 33.06% 22.12% Light Sensor 95.24% 83.33% 88.89% 95.19%
+Average (ID) 23.45% 47.57% 26.58% 24.62% 43.36% 85.64% 53.03% 44.01% 90.93% 85.64% 87.69% 91.42%
 State of Object / Location
-Faucet Open 80.00% 97.78% 88.00% 82.86% Humidity 93.62% 97.78% 95.65% 94.29%
-Stove Open 79.22% 87.14% 82.99% 77.27% Temperature 96.83% 87.14% 91.73% 90.00%
-Average (State) 79.61% 92.46% 85.50% 80.06% 95.22% 92.46% 93.69% 92.14%
+Faucet Open 72.83% 74.44% 73.63% 65.71% 80.00% 97.78% 88.00% 82.86% Humidity 93.62% 97.78% 95.65% 94.29%
+Stove Open 69.49% 58.57% 63.57% 57.27% 79.22% 87.14% 82.99% 77.27% Temperature 96.83% 87.14% 91.73% 90.00%
+Average (State) 71.16% 66.51% 68.60% 61.49% 79.61% 92.46% 85.50% 80.06% 95.22% 92.46% 93.69% 92.14%
 Surveillance
-Food Burning 50.91% 70.00% 58.95% 51.25% PM 93.33% 70.00% 80.00% 82.50%
-Chemical Spill 25.40% 80.00% 38.55% 43.33% Gas (Alcohol) 88.89% 80.00% 84.21% 93.33%
-Average (Sur.) 38.15% 75.00% 48.75% 47.29% 91.11% 75.00% 82.11% 87.91%
-Average (all) 46.54% 83.17% 55.70% 48.99% 91.93% 84.79% 87.78% 90.80%
-Table 2. Summary of end-to-end performance between RASP and camera-only for all sensing tasks.
+Food Burning 44.64% 62.50% 52.08% 42.50% 50.91% 70.00% 58.95% 51.25% PM 93.33% 70.00% 80.00% 82.50%
+Chemical Spill 18.03% 55.00% 27.16% 34.44% 25.40% 80.00% 38.55% 43.33% Gas (Alcohol) 88.89% 80.00% 84.21% 93.33%
+Average (Sur.) 31.34% 58.75% 39.62% 38.47% 38.15% 75.00% 48.75% 47.29% 91.11% 75.00% 82.11% 87.91%
+Average (all) 31.18% 51.74% 34.46% 32.17% 46.54% 83.17% 55.70% 48.99% 91.93% 84.79% 87.78% 90.80%
+Table 2. Summary of end-to-end performance between FlexiFly and camera-only for all sensing tasks.
 We ran 2-3 tasks in each of the categories of tasks, as we discuss next. For each task, we issued 70 different trials. The scenarios are described in more detail next.
@@
 In this surveillance task, users ask “let me know when my food is burning”.
-For safety, we simulate burning food by boiling water in a kettle, releasing water vapors.
+We simulate burning food by boiling water.
 The system will then actuate the drone attached with a particulate matter sensor.
@@
 The system will load an actuation module with rat poison pellets (simulated with small snacks) and the drone will dispense them at the specified location.
-For non-actuation tasks, Figure 11 shows a breakdown of the visual-language model performance in identifying the locations to send the drone versus the segmentation clustering method (Section 4).
-We see that the ARCK-Means clustering method yields the highest recall, across all scenarios, due to improvements in maintaining a good aspect ratio and object grouping over other clustering methods.
-Figure 12 highlights improvements that RASP brings compared to a purely static camera-based system across all sensing tasks, especially with respect to false positives.
+7.2. Results and Analysis
+We compare FlexiFly against two baselines: camera + LLaVA setup from our preliminary study (Camera Baseline) and augmenting the camera baseline with ARCK-Means only (Camera + ARCK-Means).
+For non-actuation tasks, Figure 9 shows a breakdown of the visual-language model performance in identifying the locations to send the drone versus the clustering method (Section 4).
+We see that the ARCK-Means clustering method yields the highest recall, across all scenarios, due to improvements in maintaining common aspect ratios and whole objects over other methods.
+Figure 10 highlights improvements that FlexiFly brings compared to a purely static camera-based system across all sensing tasks.
 A successful or “accurate” trial in this context means that the system was able to correctly identify the correct object or location (object/location ID task), correctly identify the state of the object or location (object/location state task), or correctly identify when a targeted event occurs (surveillance task); any additional points of interest identified are counted as incorrect identifications (false positives).
-For the non-RASP solution, we run the entire SAM, LLaVA, DINO pipeline, but do not actuate the drone.
 Because we are often looking for small item(s) and locations in a large scene, the visual-language model pipeline often identifies multiple points of interest (e.g., DINO draws multiple bounding boxes and locations).
-Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high and the precision becomes low with additional locations identified.
-However, adding in the RASP-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity).
+Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high, and the precision becomes low with additional locations identified.
+However, adding in the FlexiFly-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity).
 This both reduces false positives and improves overall accuracy.
 Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall).
-In total, integrating RASP with FMs improved the task success rate by $85\%$.
+In total, integrating FlexiFly with FMs improved the task success rate by $85\%$.
 Scenario # of user prompts # of executions Execution
@@
 Table 3. System performance metrics across benchmarked commands.
-For the two actuation tasks, we observed an average offset of the drone, from where it was supposed to travel to drop its payload, of 9.1cm and 10.3cm for the “poison” and “medicine” tasks, respectively.
+For the two actuation tasks, we observed a median offset of the drone, from where it was supposed to travel to drop its payload, of 9.1cm and 10.3cm for the “poison” and “medicine” tasks, respectively.
 The offset is on orders of centimeters, meaning the system was able to effectively deliver items to the proper location in most cases.
@@
 7.3. Reconfiguring Mid-Mission
-Category Success P1 Success P2 Success Total
-T1: Actuate + Actuate 20/20 = 1.0 18/20 = 0.90 18/20 = 0.90
-T2: Sense + Actuate 17/20 = 0.85 20/20 = 1.0 22/25 = 0.85
-T3: Sense + Sense 19/20 = 0.95 18/19 = 0.95 18/20 = 0.90
-Table 4. Summary of success rate for tasks that require reconfiguring the drone mid-mission, broken down by success rate of the first leg (P1) and second leg (P2).
 Figure 11. Example of identified locations and measurements in our multi-step “sense + sense” task.
-The previous section demonstrated how RASP could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting.
-Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy multi-layered tasks.
-Common tasks that require RASP to reconfigure the drone mid-mission come in three different flavors: 1) actuate + actuate: multiple individual actuation tasks aggregated into one task (e.g., “bring me my medicine AND a snack”), 2) sense + sense: a task that involves multiple sensing modalities (e.g., “find me the coolest (temperature) place to sit out of the sunlight (light)”, 3) sense + actuate: tasks that involve performing actuation in response to sensing (e.g., “place rat poison (actuation) in dark areas (sensing)”).
+The previous section demonstrated how FlexiFly could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting.
+Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how FlexiFly could be easily reconfigured mid-mission to satisfy multi-layered tasks.
+Common tasks that require FlexiFly to reconfigure the drone mid-mission come in three different flavors: 1) actuate + actuate: multiple individual actuation tasks aggregated into one task (e.g., “bring me my medicine AND a snack”), 2) sense + sense: a task that involves multiple sensing modalities (e.g., “find me the coolest (temperature) place to sit out of the sunlight (light)”, 3) sense + actuate: tasks that involve performing actuation in response to sensing (e.g., “place rat poison (actuation) in dark areas (sensing)”).
 Table 4 summarizes our results, where we show the success rate of the first task (P1), the second task (P2), and the aggregate.
@@
 There were three times that a location away from the kettle was chosen (P1).
-In these instances, our image segmentation approach cut out the area we placed the kettle, so the failure point was from the camera rather than RASP.
+In these instances, our image segmentation approach cut out the area we placed the kettle, so the failure point was from the camera rather than FlexiFly.
 Improving sensing with foundational models in future work is key to realizing a robust version of this end-to-end system.
@@
 Here, we placed the kettle at location three.
-5.3. In-the-Wild Deployment
+Category Success P1 Success P2 Success Total
+T1: Actuate + Actuate 20/20 = 1.0 18/20 = 0.90 18/20 = 0.90
+T2: Sense + Actuate 17/20 = 0.85 20/20 = 1.0 22/25 = 0.85
+T3: Sense + Sense 19/20 = 0.95 18/19 = 0.95 18/20 = 0.90
+Table 4. Summary of success rate for tasks that require reconfiguring the drone mid-mission, broken down by success rate of the first leg (P1) and second leg (P2).
 Category # Prompts # Exec Exec Time (s) Success
@@
 Number of executions is the total number of tasks issued during the deployment period.
+7.4. In-the-Wild Deployment
 After benchmarking several tasks per category, we allowed people who occupied this office space (Figure 8e) to freely use the system over the course of $5$ days.
@@
 Although there are still improvements needed to realize a truly autonomous drone-based personal assistant, all users were positively receptive to this system and could see its value.
-6. Discussion
-The focus of this work is to demonstrate how a novel autonomous reconfigurable sensing and actuation drone platform is a critical step towards FMs, LLMs, and AGI that can interact, respond, and actuate our physical environments seamlessly, much like what has been achieved in the digital domain.
-To progress further, there are several directions that warrant further attention, which we discuss next.
-Robust drone path planning, navigation, and search for indoor environments.
-As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented that could not capture and detect small objects of interest in a cluttered home/indoor setting.
-More work into exploring, detecting, and navigating to these small objects of interest is needed.
-Penetrative AI, foundation models and leveraging diverse sensors in the environment.
-In this work, we leveraged cameras in conjunction with language to understand and interpret the environment.
-However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundation models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing FMs and RASP to more scenarios.
+8. Discussion and Future Work
+Usability
+The deployment of autonomous drones in indoor environments raises critical usability challenges.
+A primary concern is noise disruption from drone propellers, which could be mitigated through: 1) flight path optimization that maintains higher operating altitudes when possible, reducing perceived ground-level noise (Adlakha et al., 2023); 2) implementation of low-noise propeller designs that recent aerodynamics research suggests could reduce noise by around 5dB (Sebastian and Strem, 2020; Jansen, 2024); and 3) context-aware navigation that avoids occupied areas during noise-sensitive periods.
+Beyond noise, user interaction with the system requires streamlining.
+We envision developing natural interaction paradigms including gesture control for intuitive drone guidance (Abtahi et al., 2017), and an augmented reality interface for visualizing drone intentions and planned paths (Mourtzis et al., 2022; Konstantoudakis et al., 2022).
+Privacy
+Privacy is a critical concern for camera networks and camera-equipped drones in indoor spaces.
+In this work, camera feeds are transmitted locally and processed on a local server for scene understanding, vision-language grounding and drone navigation.
+The need for server can be bypassed as more efficient, compact foundation models and powerful edge computers emerge, as well as leveraging compression techniques such as quantization and distillation to create dedicated smaller models (Sharshar et al., 2025; Cai et al., 2024) for each vision task mentioned above while preserving the generalizibility.
+Extending to Diverse Environments
+While we We demonstrated the effectiveness of our implementation in several indoor environments, extending FlexiFly to new settings presents unique challenges and opportunities.
+Flexifly can be adapted to environments with existing camera infrastructure as long as the relative position of the cameras are known; this can be achieved in a self-supervised manner by leveraging recent advances in camera self-localization and calibration techniques (Liu et al., 2022), potentially enhanced by fine-tuning vision-language models for specific deployment contexts.
+However, we acknowledge significant limitations in environments where camera deployment is impractical or restricted.
+For spaces primarily monitored by non-visual sensors (e.g., RF, vibration, or acoustic sensors), new methodologies beyond ARCK-Means and our vision pipeline must be developed for localizing areas of interest and decision-making.
+We envision that future FMs and penetrative AI (Xu et al., 2024) could help process diverse sensor modalities spread throughout environments, aided by novel segmentation approaches for non-visual spatial data to enable accurate drone navigation.
 9. Conclusion
-We propose RASP, a modular and reconfigurable drone sensing and actuation platform that allows FMs and LLMs to dynamically to swap sensing and actuation capabilities on-the-fly.
-We demonstrate through a real deployment of a personal assistant application that RASP can improve the successful completion of complex tasks throughout our physical spaces by up to $85\%$.
-This drastic improvement stems from RASP’s ability to provide localized sensing and actuation in a large space even without large deployments of sensors and devices, as is commonly seen today.
-RASP is a critical step towards FMs and LLMs that can naturally interact and actuate the physical environment, just as they have shown in many applications in the digital domain.
+Our work studies the possibility of LLMs and FMs as a general intelligence for physical spaces.
+We identify that FMs analyzing sensing data monitoring a large space have difficulty identifying localized events that occur in small areas.
+As such, we propose novel segmentation methods and drones with reconfigurable sensing and actuation that enable FMs to identify and “zoom in” to analyze targeted areas with higher resolution.
+We demonstrate through a real deployment of a personal assistant application that FlexiFly can improve the successful completion of complex tasks throughout our physical spaces by up to $85\%$.
+FlexiFly is a critical step towards FMs and LLMs that can naturally interact and actuate the physical environment, just as they have shown in many applications in the digital domain.
+Acknowledgements.
+This research was partially supported by COGNISENSE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA, as well as the National Science Foundation under Grant Number CNS-1943396.
+The views and conclusions contained here are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Columbia University, NSF, SRC, DARPA, or the U.S.
+Government or any of its agencies.